### v1

Abstract
In this paper, for Markov decision processes (MDPs) with unbounded state spaces we present refined upper bounds presented in [Kara et. al. JMLR’23] on finite model approximation errors via optimizing the quantizers used for finite model approximations. We also consider implications on quantizer design for quantized Q-learning and empirical model learning, and the performance of policies obtained via Q-learning where the quantized state is treated as the state itself. We highlight the distinctions between planning, where approximating MDPs can be independently designed, and learning (either via Q-learning or empirical model learning), where approximating MDPs are restricted to be defined by invariant measures of Markov chains under exploration policies, leading to significant subtleties on quantizer design performance, even though asymptotic near optimality can be established under both setups. In particular, under Lyapunov growth conditions, we obtain explicit upper bounds which decay to zero as the number of bins approaches infinity.
keywords:
Reinforcement learning, Quantizer design, MDP
1 Introduction
It has been recently shown that one can obtain finite approximations via state and action quantization for Markov decision processes (MDPs) with uncountable Polish spaces, as well as run both empirical model learning and Q-learning to arrive at near optimal solutions (see.e.g. [24, 27]). Such studies have primarily focused on the case with compact spaces and uniform quantization and with only asymptotic convergence results for the case with non-compact spaces. To this end, the goal of this paper is to design quantizers in such a finite model approximation and learning framework for continuous space MDPs. Notably, this paper develops a mathematical framework and formulation paving the way for non-uniform quantizer design and associated performance bounds, which are then also applicable to MDPs with non-compact spaces.
1.1 Related Literature
For stochastic control problems with continuous state and action spaces, approximations are inevitable for computational methods. A common approach in reinforcement learning is function approximation, where the value function (Q-function) of the control problem is approximated using a parametrized family of functions. Convergence of policy evaluation methods are known under linear function approximation where the parametrized family of functions are formed by the span of finitely many basis functions [33]. However, learning optimal Q-functions with linear function approximation is known to be unstable in general [2] except in special cases. For general linear function approximation, [20] has shown that under certain a class of exploration policies, optimal Q-value learning with linear function approximation remains bounded. The special cases where the convergence can be guaranteed include when (i) the exploration policy is already close to the greedy policy of the learning iterations [18], and (ii) the stage-wise cost function, the transition model, and thus the optimal Q-functions are perfectly represented by the basis functions, i.e. they belong to the span of the basis functions [10, 23]. These assumptions, however, are restrictive in general, since it is unrealistic to assume near-optimal exploration or perfect linear representability.
A particularly powerful, though computationally demanding, special case is when the basis functions are indicator functions of quantization bins, which form an orthonormal basis, and the convergence analysis can mathematically be justified. In particular, [16] showed that the Q-values learned under quantization based learning correspond to an auxiliary finite control problem, which is a finite approximation of the original control problem with a particular weight measure on the quantization bins. This observation has an important consequence that the model-based approaches using state space quantization and the quantized Q-learning coincide and can be used interchangeably, even under very mild continuity conditions on the transition kernels [27].
For quantization based learning methods while often convergence is studied, error analysis regarding the limit of the stochastic iterates is typically not studied in general or an error analysis is not provided at all. Early studies in this direction include [31, 8]. [32] generalized these by the use of Q function interpolators that are sufficiently regular (defined by non-expansiveness) in their parametric representation and established both convergence and optimality properties. Error analysis of the learned value functions with respect to the true value functions under quantization methods is studied in [29], where the authors established finite-sample guarantees for quantized Q-learning with nearest-neighbor mappings, assuming transition models admit continuous densities with respect to the Lebesgue measure.
In [27, 16, 17], it was shown that the approximate value function, as well as the performance of the resulting policies, are nearly optimal under weakly continuous (or weak Feller) models, arguably the weakest assumption ensuring existence of optimal solutions and consistency of aproximations.
Another direction of research with provable loss guarantees is kernel-based methods [22, 21, 3, 6, 35, 7] where the cost and the transition models are estimated via empirical kernel regression or otherwise simulations are used to update the Bellman optimality equations. One can then find the control policies based on the learned models using model-based approaches.
Most prior work focuses on approximation of control problems with compact and bounded spaces. There is a limited number of provable guarantees for the approximation of control problems with unbounded state spaces. In [27] asymptotic near optimality of the quantization based approximations was established for non-compact state spaces as the quantization rate grows to infinity. However, there is no convergence rate guarantees for this approach in terms of expected quantization error on the state space and learning was not studied in this context.
Moreover, previous work generally assumes a fixed quantization scheme, without focusing on the design of the quantizers. In this design direction, [30] studied adaptive quantization methods for compact spaces under Wasserstein continuous transitions, proposing progressively refined partitions based on ‘relevance’ of the partition sets. The resulting quantization scheme is not optimal in general, as refinement is localized rather than globally optimized.
1.2 Model and Cost Criteria
Let $\mathds{X}\subset\mathds{R}^{n}$ be a Borel set in which the elements of a controlled Markov chain $\{X_{t},\,t\in\mathbb{Z}_{+}\}$ take values for some $n<\infty$. Here and throughout the paper, $\mathbb{Z}_{+}$ denotes the set of non-negative integers and $\mathds{N}$ denotes the set of positive integers. Let $\mathds{U}$, the action space, be a compact Borel subset of some Euclidean space, from which the sequence of control action variables $\{U_{t},\,t\in\mathbb{Z}_{+}\}$ take values.
The $\{U_{t},\,t\in\mathbb{Z}_{+}\}$, are generated via admissible control policies: An admissible policy $\gamma$ is a sequence of control functions $\{\gamma_{t},\,t\in\mathbb{Z}_{+}\}$ such that $\gamma_{t}$ is measurable on the $\sigma$-algebra generated by the information variables
$I_{t}=\{X_{0},\ldots,X_{t},U_{0},\ldots,U_{t-1}\},\quad t\in\mathds{N},\quad\quad I_{0}=\{X_{0}\},$
where
$U_{t}=\gamma_{t}(I_{t}),\quad t\in\mathbb{Z}_{+},$ (1)
are the $\mathds{U}$-valued control actions. We define $\Gamma$ to be the set of all such admissible policies.
The joint distribution of the state and control processes is then completely determined by (1), the initial probability measure of $X_{0}$, and the following relationship:
$\mathop{\rm Pr}\biggl(X_{t}\in B\,\bigg|\,(X,U)_{[0,t-1]}=(x,u)_{[0,t-1]}\biggr)=\int_{B}\mathcal{T}(dx_{t}|x_{t-1},u_{t-1}),B\in\mathcal{B}(\mathds{X}),t\in\mathds{N},$ (2)
where $\mathcal{T}(\cdot|x,u)$ is a stochastic kernel (that is, a regular conditional probability measure) from $\mathds{X}\times\mathds{U}$ to $\mathds{X}$, $\mathcal{B}(\mathds{X})$ is the Borel $\sigma$-algebra of $\mathds{X}$, and $(X,U)_{[0,t-1]}$ is the set of state-action pairs up until t1t-1.
We will be interested in the following performance criteria: The first one is the infinite-horizon discounted expected cost
$J_{\beta}(x_{0},\gamma)=E_{x_{0}}^{{\mathcal{T}},\gamma}\left[\sum_{t=0}^{\infty}\beta^{t}c(X_{t},U_{t})\right]$ (3)
where $0<\beta<1$ is the discount factor, $c:\mathds{X}\times\mathds{U}\to\mathds{R}$ is the stage-wise continuous and bounded cost function, and $E_{x_{0}}^{{\mathcal{T}},\gamma}$ denotes the expectation with initial state $x_{0}$ and transition kernel $\mathcal{T}$ under policy $\gamma$. Furthermore, for any initial state $X_{0}=x_{0}$, the optimal value function is defined by
$J_{\beta}^{*}(x_{0})$ $=\inf_{\gamma\in\Gamma}J_{\beta}(x_{0},\gamma).$
The second objective is the infinite horizon average cost criterion
$J_{avg}^{*}(x):=\inf_{\gamma}J_{avg}(x,\gamma)=\inf_{\gamma\in\Gamma}\limsup_{T\to\infty}{1\over T}E^{\gamma}_{x}[\sum_{t=0}^{T-1}c(x_{t},u_{t})].$ (4)
1.3 Contributions
In this paper, we make the following contributions:
  • (i)
    In Theorem 2, we refine and computationally improve the upper bounds given in [16] which involve admissible policies to ones that only involve stationary policies. This facilitates an analysis involving occupation measures as well as invariant measures for discounted and ergodic cost criteria, respectively; which is then utilized later in the paper.
  • (ii)
    In Theorems 3 9, we derive bounds in terms of occupation measures, to represent the loss in terms of occupation and invariant measures, respectively for the discounted and average cost criteria. In Corollary 6, we optimize the quantization design by choosing the representative points as medians (in the $\ell_{1}$ sense) with respect to the occupation measures and arrive at an explicit error bound. We thus provide a convergence rate analysis of quantization based approximations which is convenient for stochastic analysis and which is also applicable for non-compact state spaces.
  • (iii)
    In Theorems 7 and 8 under Foster-Lyapunov conditions, we derive explicit error bounds which decay to zero as the quantization gets finer and show better performance than uniform quantization in MDPs with non-compact state spaces under discounted cost criteria. We extend our analysis to the average cost criterion by obtaining explicit error bounds in Theorem 11.
  • (iv)
    In Theorem 14, we extend our analysis on quantizer design to quantized Q-learning and empirical model learning (which are equivalent in performance). We obtain the error bounds which depend, unlike the planning problem above, on the invariant measure induced by the exploration policy in Theorem 13. For non-compact spaces, in Theorem 14, we show that the approximation error diminishes as quantization becomes finer under Foster–Lyapunov conditions, despite constraints on the dependence of the weighting measures on the exploration policy and quantization bins, and we derive explicit error bounds.
2 Refined Error Bounds on Finite Model Approximations
2.1 Approximate Model Construction
We start with the the approach introduced in [25, 28], where we partition the state space $\mathbb{X}$ into $M$ disjoint subsets $\{B_{i}\}_{i=1}^{M}$, such that $\bigcup_{i=1}^{M}B_{i}=\mathbb{X}$ and $B_{i}\cap B_{j}=\emptyset$ for $i\neq j$. For each subset $B_{i}$, we select a representative state $y_{i}\in B_{i}$. The finite set $\mathbb{Y}=\{y_{1},y_{2},\dots,y_{M}\}$ serves as the quantized state space. The quantizer mapping $q:\mathbb{X}\to\mathbb{Y}$ is defined by
$q(x)=y_{i}\quad\text{if }x\in B_{i}.$
We introduce a probability measure $\pi\in\mathcal{P}(\mathbb{X})$ over $\mathbb{X}$, ensuring that $\pi(B_{i})>0$ for each $B_{i}$. This measure allows us to define normalized measures for each quantization bin $B_{i}$:
$\hat{\pi}_{y_{i}}(A)=\frac{\pi(A)}{\pi(B_{i})},\quad\forall A\subseteq B_{i},\quad\forall i\in\{1,\ldots,M\}.$
Using these normalized measures, we define the stage-wise cost and transition kernel for the finite-state MDP:
$C^{*}(y_{i},u)=\int_{B_{i}}c(x,u)\hat{\pi}_{y_{i}}(dx),$
$P^{*}(y_{j}|y_{i},u)=\int_{B_{i}}\mathcal{T}(B_{j}|x,u)\hat{\pi}_{y_{i}}(dx).$
The finite-state value function $\hat{J}_{\beta}:\mathds{Y}\rightarrow\mathbb{R}$ satisfies the dynamic programming equation:
$\hat{J}_{\beta}(y)=\inf_{u\in\mathbb{U}}\left\{C^{*}(y,u)+\beta\sum_{z\in\mathbb{Y}}\hat{J}_{\beta}(z)P^{*}(z|y,u)\right\}.$
We extend $\hat{J}_{\beta}$ to $\mathbb{X}$ by setting $\hat{J}_{\beta}(x)=\hat{J}_{\beta}(q(x))$ for all $x\in\mathbb{X}$.
Under certain regularity conditions, we will see that the quantization error can be efficiently bounded by the loss function $L:\mathbb{X}\rightarrow\mathbb{R}$:
$L(x)=\int_{B_{i}}\|x-x^{\prime}\|_{1}\hat{\pi}_{y_{i}}(dx^{\prime}),\quad\forall x\in B_{i}$ (5)
where $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm in $\mathds{R}^{n}$.
Remark 2.1.
We note that the loss function (5) above is also often called the potential function of the measure $\hat{\pi}_{y_{i}}$ evaluated at $x$, in the mathematical statistics and probability theory literatures, see e.g. [4].
2.2 Refined Error Bounds for the Discounted Cost Criterion
To derive error bounds, we make the following assumptions:
Assumption 2.1.
The MDP $(\mathds{X},\mathds{U},\mathcal{T},c)$ satisfies the following:
  • (i)
    The action space $\mathds{U}$ is compact.
  • (ii)
    The stage-wise cost function $c$ is nonnegative, bounded and continuous on both $\mathds{X}$ and $\mathds{U}$.
  • (iii)
    The kernel $\cal T$} is weak Feller, that is, for every $g\in C_{b}(\mathds{X})$ (continuous and bounded function), the map
    $\mathds{X}\times\mathds{U}\ni(x,u)\mapsto\int g(x_{1}){\cal T}(dx_{1}|x_{0}=x,u_{0}=u)\in\mathbb{R}$
    is continuous.
The above ensures that optimal policies exist. Furthermore, by [24, Lemma 3.19] and [24, Theorem 3.16] (see also [26]), any MDP with a weakly continuous transition probability kernel can be approximated by an MDP with finite action spaces. Accordingly, in the sequel, we assume that $\mathds{U}$ is finite.
Remark 2.2.
[…]
Assumption 2.2.
The transition kernel and the stage-wise cost function satisfies the following:
  • (i)
    $c(x,u)$ is Lipschitz continuous in $x$. There exists a constant $\alpha_{c}>0$ such that
    $|c(x,u)-c(x^{\prime},u)|\leq\alpha_{c}\|x-x^{\prime}\|_{1},\quad\forall x,x^{\prime}\in\mathds{X},\forall u\in\mathds{U}.$
  • (ii)
    $\mathcal{T}(\cdot|x,u)$ is Lipschitz continuous in $x$ under the total variation distance. There exists a constant $\alpha_{T}>0$ such that
    $\|\mathcal{T}(\cdot|x,u)-\mathcal{T}(\cdot|x^{\prime},u)\|_{\text{TV}}\leq\alpha_{T}\|x-x^{\prime}\|_{1},\quad\forall x,x^{\prime}\in\mathds{X},\forall u\in\mathds{U}.$
Under these assumptions, we first recall the following theorem:
Theorem 1 (Kara et al., 2023, Theorem 3 [16]).
Under Assumptions 2.1 and 2.2, for any initial state $x_{0}\in\mathds{X}$, the error between the optimal value function $J^{*}_{\beta}(x_{0})$ and the approximate value function $\hat{J}_{\beta}(x_{0})$ satisfies:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sum_{t=0}^{\infty}\beta^{t}\sup_{\gamma\in\Gamma}\mathbb{E}_{x_{0}}^{\gamma}[L(X_{t})],$
where $\Gamma$ is the set of admissible policies, and $L(X_{t})$ is the loss function defined in (5).
The bound in Theorem 1 involves a supremum over all admissible policies $\Gamma$, which can be difficult to compute. In the following, we refine the bound by restricting the supremum to stationary policies, which will turn out to be consequential in our analysis to follow as this will allow the bounds to be computed in terms of occupation measures or invariant measures.
Theorem 2.
Under Assumptions 2.1 and 2.2, for any initial state $x_{0}\in\mathds{X}$, the error satisfies:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\mathbb{E}_{x_{0}}^{\gamma_{s}}\left[\sum_{t=0}^{\infty}\beta^{t}L(X_{t})\right],$
where $\gamma_{s}$ is the policy that achieves the supremum for $\sup_{u}\int|\hat{J}_{\beta}(x_{1})-{J}^{*}_{\beta}(x_{1})|\mathcal{T}(dx_{1}|x,u)$.
In particular, we have that
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\mathbb{E}_{x_{0}}^{\gamma_{s}}\left[\sum_{t=0}^{\infty}\beta^{t}L(X_{t})\right],$ (6)
where $\Gamma_{s}$ is the set of stationary policies.
Proof.
The proof is in Appendix A
We introduce the discounted occupation measure. For any measurable set $D\in\mathcal{B}(\mathbb{X}\!\times\!\mathbb{U})$ we define
$\nu_{x_{0}}^{\gamma_{s}}(D)=\sum_{t=0}^{\infty}\beta^{t}\,\mathbb{E}_{x_{0}}^{\gamma_{s}}\!\bigl[\mathbf{1}_{D}(X_{t},U_{t})\bigr]=\sum_{t=0}^{\infty}\beta^{t}\,\mathbb{P}_{x_{0}}^{\gamma_{s}}\!\bigl((X_{t},U_{t})\in D\bigr),$
where the probability measure $\mathbb{P}_{x_{0}}^{\gamma_{s}}$ over the state and action process is defined by the initial condition $x_{0}$, the stationary policy $\gamma_{s}$, and the transition kernel $\mathcal{T}$.
For a product set $A\times\mathds{U}$ with $A\subseteq\mathds{X}$ measurable we obtain
$\nu_{x_{0}}^{\gamma_{s}}(A\times\mathds{U})=\sum_{t=0}^{\infty}\beta^{t}\,\mathbb{P}_{x_{0}}^{\gamma_{s}}\bigl(X_{t}\in A\bigr)=\sum_{t=0}^{\infty}\beta^{t}\,\mu_{t}^{\gamma_{s}}(A)=:\frac{1}{1-\beta}\,\mu_{\beta}^{\gamma_{s}}(A),$
where $\mu_{\beta}^{\gamma_{s}}$ is a probability measure on $(\mathds{X},\mathcal{B}(\mathds{X}))$ obtained by
$\mu_{\beta}^{\gamma_{s}}(A):=(1-\beta)\sum_{t=0}^{\infty}\beta^{t}\,\mu_{t}^{\gamma_{s}}(A).$
Theorem 3.
Under Assumption 2.1 and Assumption 2.2, and given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}\in\mathds{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|$ $\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{1}{1-\beta}\sup_{\gamma_{s}\in\Gamma_{s}}\int_{\mathds{X}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)$
where $\Gamma_{s}$ represents the set of stationary policies. The term $\mu_{\beta}^{\gamma_{s}}(dx)$ denotes the normalized discounted occupation measure over the state space.
Proof.
The proof is in Appendix B. ∎
To minimize the error bound obtained, we optimize the quantizer by choosing the centroids of the quantization bins properly. The following lemma will be useful in choosing the centroids.
Lemma 4.
[5] Let $\mathds{X}\subseteq\mathds{R}$, and let $X$ be a random variable with distribution $\mu$ over a quantization bin $B_{i}$. Suppose that $\mathbb{E}(|X|)<\infty$. Then, the value $y_{i}$ that minimizes the expected distortion $\mathbb{E}[|X-y_{i}|]$ is the median of $X$ with respect to the probability measure $\mu$ over $B_{i}$ .
For a proof, see [5].
According to Lemma 4, we establish that for a given quantization bin $B_{i}$, the expected distortion is minimized when the representative point $y_{i}$ is chosen as the median of $B_{i}$ with respect to the probability measure $\mu$ over $B_{i}$. This turns the normalized measure $\hat{\pi}_{y_{i}}$ into the Dirac measure $\delta_{y_{i}}$ concentrated at $y_{i}$. This result can be also applied to higher dimensional state spaces in $\mathbb{R}^{n}$ according to the following theorem.
The result can be extended to higher dimensions when considering the $\ell_{1}$ norm in $\mathbb{R}^{n}$.
Lemma 5.
Let $X=(X_{1},X_{2},\dots,X_{n})$ be a random vector with finite expected absolute deviations in each coordinate. Then, the point $y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*})$ that minimizes $\mathbb{E}[\|X-y_{i}\|_{1}]$ is obtained by choosing each $y_{i,k}^{*}$ to be a median of the marginal distribution of $X_{k}$ over the bin $B_{i}$, for $k=1,2,\dots,n$.
The proof of the above is immediate: Let $X$ be an n -dimensional random vector, and $y_{i}\in\mathbb{R}^{n}$ . The expected $\ell_{1}$ distortion is:
$\mathbb{E}[\|X-y_{i}\|_{1}]=\int_{B_{i}}\|x-y_{i}\|_{1}\mu(dx)=\sum_{k=1}^{n}\int_{B_{i}}|x_{k}-y_{i,k}|\mu_{k}(dx),$ (7)
where $\mu_{k}$ is the marginal distribution of $X_{k}$ over the bin $B_{i}$. Since the sum of convex functions is convex, $\mathbb{E}[\|X-y_{i}\|_{1}]$ is convex in $y_{i}$ . Each component $y_{i,k}$ appears only in the term involving $x_{k}$, so we can minimize each term separately. By the same reasoning as in the scalar case, as in Lemma 4, $\int_{B_{i}}|x_{k}-y_{i,k}|\mu_{k}(dx)$ is minimized when $y_{i,k}$ is a median of $X_{k}$ with respect to the probablity measure $\mu_{k}$ over $B_{i}$. Therefore, the point $y^{*}_{i}$ that minimizes the expected $\ell_{1}$ distortion is found as:
$y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*}),\quad\text{for}\quad k=1,2,\dots,n$ (8)
where each $y_{i,k}^{*}$ is the median of the quantization bin $B_{i}$ with respect to the marginal distribution $\mu_{k}$.
Corollary 6 (to Theorem 3 and Theorem 5).
Under Assumption 2.1 and Assumption 2.2, and given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}\in\mathds{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\|x-y_{i}\|_{1}\mu_{\beta}^{\gamma_{s}}(dx),$
where $\Gamma_{s}$ represents the set of stationary policies, and $y_{i}$ is the median of the quantization bin $B_{i}$ under the measure $\mu_{\beta}^{\gamma_{s}}(dx)$ where $\mu_{\beta}^{\gamma_{s}}(dx)$ denotes the normalized discounted occupation measure over the state space.
Our following result illustrates how the expected loss during the application of a quantization can be bounded by the use of a Lyapunov function for non-compact state spaces.
Theorem 7.
Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $\alpha>0$ and $V:\mathds{X}\to[0,\infty)$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
$\mathbb{E}^{\gamma_{s}}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-\alpha V(x)+b,\quad x\in\mathds{X},u\in\mathds{U}$ (9)
for all $\gamma_{s}\in\Gamma_{s}$ where $V(x)=|x|^{m}$, for some $m>1$.
Then, under Assumption 2.1 and Assumption 2.2, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathds{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{3C^{1/m}}{(M^{1-1/m})(1-\beta)},$
where $C$ is a constant, which depends on the initial state $x_{0}$, defined as:
$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
Proof.
The proof is in Appendix C. ∎
We note that the existence discussion is constructive and the proof of Theorem 7 utilizes an explicit quantizer which attains the bound presented. Observe that as $M\to\infty$, the error converges to zero.
The next result extends Theorem 7 to an nn–dimensional state space $\mathds{X}\subseteq\mathbb{R}^{n}$ (with the $\ell_{1}$–norm).
Theorem 8.
Let $b\geq 0$, $\alpha>0$, and define the Lyapunov function $V(x)=\|x\|_{1}^{m}$ with exponent $m>n$. Assume the controlled process $\{X_{t}\}$ satisfies the drift condition
$\mathbb{E}^{\gamma_{s}}\bigl[V(X_{t+1})\mid X_{t}=x,U_{t}=u\bigr]\leq V(x)-\alpha V(x)+b,\qquad x\in\mathds{X},u\in\mathds{U}$
for all $\gamma^{s}\in\Gamma_{s}$. Under Assumption 2.1 and Assumption 2.2, let $M$ be the total number of hyper-cubic bins in the uniform quantizer. Then, for every initial state $x_{0}\in\mathds{X}$ we have that,
$\bigl|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\bigr|\leq\Bigl(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\Bigr)\frac{(2n+1)C^{1/m}}{M^{\frac{1}{n}-\frac{1}{m}}(1-\beta)},$ (10)
where
$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
Proof.
The proof is in Appendix D. ∎
The explicit quantizer is presented in the proof of Theorem 8.
2.3 Refined Error Bounds for the Average Cost Criterion
In this section, we extend the analysis to the average cost criterion in Markov Decision Processes (MDPs). We focus on the long-run average cost criterion, defined as:
$J_{\text{avg}}(x_{0},\gamma)=\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}^{\gamma}_{x_{0}}\left[\sum_{t=0}^{T-1}c(X_{t},U_{t})\right],$
where $\gamma=\{\gamma_{t}\}_{t=0}^{\infty}$ is the policy, and $J_{\text{avg}}(x_{0})$ represents the average cost starting from state $x_{0}$.
We now define the average cost problem as follows:
$J_{\text{avg}}^{*}(x):=\inf_{\gamma}J_{\text{avg}}(x,\gamma)=\inf_{\gamma\in\Gamma_{A}}\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{x}^{\gamma}\left[\sum_{t=0}^{T-1}c(X_{t},U_{t})\right]$
where $\Gamma_{A}$ represents the set of all admissible policies.
Assumption 2.3 (Minorization Condition).
There exists a non-trivial positive measure $\mu$ on $\mathds{X}$ such that for all $(x,u)\in\mathds{X}\times\mathds{U}$:
$\mathcal{T}(B\mid x,u)\geq\mu(B),\quad\forall B\in\mathcal{B}(\mathds{X}).$ (11)
Before the result, we introduce the Average Cost Optimality Equations (ACOEs) for the original and the finite model:
$h(x)=\inf_{u\in\mathds{U}}\left\{c(x,u)+\int_{\mathds{X}}h(x_{1})\mathcal{T}(dx_{1}|x,u)\right\}-\int_{\mathds{X}}h(x)\mu(dx)$
$\hat{h}(y)=\inf_{u\in\mathds{U}}\left\{C^{*}(y,u)+\sum_{y_{1}}\hat{h}(y_{1})P^{*}(y_{1}|y,u)\right\}-\int_{\mathds{X}}\hat{h}(q(x))\mu(dx).$
Existence of the solutions to these equations is guaranteed under Assumption 2.3. We will refer to the functions $h$ and $\hat{h}$ as the relative value functions in the following.
Theorem 9.
Under Assumptions 2.1, 2.2 and 2.3, $J_{\text{avg}}^{*}(x)$ and $\hat{J}_{\text{avg}}(x)$ are constants for any $x\in\mathds{X}$:
$j^{*}=J_{\text{avg}}^{*}(x)\quad\text{for all }x\in\mathds{X},$
$\hat{j}=\hat{J}_{\text{avg}}(x)\quad\text{for all }x\in\mathds{X}$
Moreover,
$|J_{\text{avg}}^{*}(x_{0})-\hat{J}_{\text{avg}}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\int_{\mathds{X}}L(x)\pi_{\gamma_{s}}(dx),$ (12)
where $\gamma_{s}$ is the policy that achieves the supremum for $\sup_{u}\int|h(x_{1})-\hat{h}(x_{1})|\mathcal{T}(dx_{1}|x,u)$ and where $\pi_{\gamma_{s}}$ is the invariant measure induced by policy $\gamma_{s}$.
Proof.
The proof is in Appendix E
Now, we state the following theorem, which will be useful in the proof of our next result:
Theorem 10 (Comparison Theorem).
[19, Theorem 14.2.2] Let $V:\mathds{X}\to[0,\infty)$ and $f,g:\mathds{X}\to[0,\infty)$. Let $\{x_{n}\}$ be a Markov chain on $\mathds{X}$. If the following condition is satisfied:
$\int_{\mathds{X}}P(x,dy)V(y)\leq V(x)-f(x)+g(x),\quad x\in\mathds{X},$
then, for any stopping time $\tau$ with $\mathbb{P}(\tau<\infty)=1$, it holds that
$\mathbb{E}\left[\sum_{t=0}^{\tau-1}f(x_{t})\right]\leq V(x_{0})+\mathbb{E}\left[\sum_{t=0}^{\tau-1}g(x_{t})\right].$
Proof.
See [34]. ∎
Now, we present our result under Lyapunov condition similar to the quantizer design for discounted cost criterion.
Theorem 11.
Assume $\mathds{X}\subseteq\mathbb{R}$, $f(x)=\|x\|^{m}$ with $m>1$, and suppose
$\sup_{x,u}\mathbb{E}[V(X_{t+1})\mid X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$
where $b\geq 0$ and $V:\mathds{X}\to[0,\infty)$ is a Lyapunov function.
Then, under Assumption 2.1, Assumption 2.2 and Assumption 2.3, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathds{X}$:
$|\hat{J}_{avg}(x_{0})-J_{avg}^{*}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{M^{1-1/m}}.$ (13)
Proof.
The proof is in Appendix F
As earlier, as $M\to\infty$ the error converges to zero.
3 Quantizer Design for Quantized Q-Learning and Empirical Model Learning
In this section, we study the design of quantizers for the Q-learning algorithm, and via equivalence, to empirical model learning. To this end, we first refine the error expression obtained in [14], and obtain an explicit error bound by extending the analysis presented in Section 2.2, and also design quantizers applicable for unbounded state spaces.
3.1 Quantized Q-Learning Algorithm and its Convergence
As noted in [14], quantizing the state space of a continuous MDP converts the problem into a Partially Observable Markov Decision Process (POMDP) which is then non-Markovian. In particular, with the quantization process, the controller has partial observation. Define
$Y_{t}=q(X_{t}),$
where $q:\mathbb{X}\rightarrow\mathbb{Y}$ is the quantizer, which maps the continuous state space $\mathbb{X}$ to the finite set of representative points $\mathbb{Y}=\{y_{1},y_{2},\ldots,y_{M}\}$.
The observation model in our setup is described as:
$O(y_{i}\mid x)=\mathbb{P}(Y_{t}=y_{i}\mid X_{t}=x)=\mathds{1}_{B_{i}}(x),$
where $\mathds{1}_{B_{i}}(x)$ is the indicator function. In this model, the agent has access to only the representative points $Y_{t}=y_{i}$, which indicates that the true state $X_{t}$ resides in the quantization bin $B_{i}$. In other words, the true state $X_{t}$ remains hidden from the agent. In the quantized MDP, the agent observes $Y_{t}=q(X_{t})$.
Now, consider the following Q-learning update rule for $(X_{T},U_{T})=(x,u)\in\mathbb{X}\times\mathbb{U}$:
$Q_{t+1}(q(x),u)=(1-\alpha_{t}(q(x),u))Q_{t}(q(x),u)+\alpha_{t}(q(x),u)\left(c(x,u)+\beta\min_{v\in\mathbb{U}}Q_{t}(q(X_{t+1}),v)\right),$ (14)
where, $\alpha_{t}(y_{t},u_{t})$ is the learning rate and $c(x,u)$ is the immediate cost which depends on the true state $x$ and action $u$. The quantized Q-learning algorithm is then implemented as:
Algorithm 1 Quantized Q-Learning Algorithm
Input: Initial Q-function $Q_{0}$, quantizer $q:\mathbb{X}\to\mathbb{Y}$, exploration policy $\gamma^{*}$, total iterations $L$
 Initialize counts $N(y,u)=0$ for all $(y,u)\in\mathbb{Y}\times\mathbb{U}$
for $t=0$ to L1L-1 do
  Observe the state $X_{t}$ and quantize the state according to $y_{t}=q(X_{t})$
  Select action $u_{t}$ according to the exploration policy $\gamma^{*}$
  Execute action $u_{t}$, receive cost $c(X_{t},u_{t})$, observe next state $X_{t+1}$
  Observe next quantized state $y_{t+1}=q(X_{t+1})$
  Update the count: $N(y_{t},u_{t})\leftarrow N(y_{t},u_{t})+1$
  Update the learning rate:
$\alpha_{t}(y_{t},u_{t})=\frac{1}{1+N(y_{t},u_{t})}$ (15)
  Update the Q-function:
$Q_{t+1}(y_{t},u_{t})=(1-\alpha_{t}(y_{t},u_{t}))Q_{t}(y_{t},u_{t})+\alpha_{t}(y_{t},u_{t})\left(c(X_{t},u_{t})+\beta\min_{v\in\mathbb{U}}Q_{t}(y_{t+1},v)\right)$ (16)
end for
Output: Learned Q-function $Q_{L}$
Assumption 3.1.
  1. 1.
    We define the step size $\alpha_{t}(y,u)$ as follows:
    $\alpha_{t}(y,u)=\begin{cases}0&\text{if }(Y_{t},U_{t})\neq(y,u),\\ \frac{1}{1+\sum_{k=0}^{t}\mathds{1}\{Y_{k}=y,U_{k}=u\}}&\text{otherwise}.\end{cases}$
  2. 2.
    Under the exploration policy $\gamma^{*}$, the state process $\{X_{t}\}_{t\geq 0}$ is uniquely ergodic, implying the existence of a unique invariant measure $\pi_{\gamma^{*}}$.
  3. 3.
    During the exploration phase, each observation-action pair $(y,u)$ is visited infinitely often. This ensures sufficient exploration of the state-action space.
We note that a sufficient condition for the second item above is that the state process $\{X_{t}\}_{t\geq 0}$ is positive Harris recurrent.
The following result is proved by Kara and Yüksel [11]. It states that the algorithm in (1) converges to the optimal Q-function of the approximate belief MDP under Assumption 3.1:
We have the following convergence result:
Theorem 12 (Kara et al., 2023, Theorem 9 [16]).
Under Assumption 3.1, the quantized Q-learning algorithm in 1 converges almost surely to a function $Q^{*}(y,u)$ that satisfies the following fixed-point equation for every $(y,u)\in\mathbb{Y}\times\mathbb{U}$:
$Q^{*}(y,u)=C^{*}(y,u)+\beta\sum_{y^{\prime}\in\mathbb{Y}}P^{*}(y^{\prime}|y,u)\min_{v\in\mathbb{U}}Q^{*}(y^{\prime},v),$ (17)
where:
$C^{*}(y,u)$ $=\frac{1}{\pi_{\gamma^{*}}(B_{y})}\int_{B_{y}}c(x,u)\pi_{\gamma^{*}}(dx),$ (18)
$P^{*}(y^{\prime}|y,u)$ $=\frac{1}{\pi_{\gamma^{*}}(B_{y})}\int_{B_{y}}\int_{B_{y^{\prime}}}\mathcal{T}(x^{\prime}|x,u)\,dx^{\prime}\,\pi_{\gamma^{*}}(dx),$ (19)
where, $B_{y}$ denotes the quantization bin corresponding to $y\in\mathbb{Y}$ and $\pi_{\gamma^{*}}$ is the invariant measure of the state process under the exploration policy $\gamma^{*}$:
3.2 Empirical Model Learning and Equivalence with Quantized Q-Learning
Let under the exploration policy $\gamma^{*}$ given in the quantized Q-learning algorithm in 1 give rise to the invariant probability measure $\pi_{\gamma^{*}}$. The limiting Q-function $Q^{*}(y,u)$ in the discussion above corresponds to the optimal Q-function of an approximate MDP defined over the quantized state space $\mathbb{Y}$. The effective cost $C^{*}(y,u)$ is the average cost over the bin $B_{y}$ weighted by the invariant distribution $\pi_{\gamma^{*}}$ conditioned on bin $B_{y}$:
$C^{*}(y,u)=\mathbb{E}_{x\sim\pi_{\gamma^{*}}|x\in B_{y}}[c(x,u)]=\int_{B_{y}}\frac{\pi_{\gamma^{*}}(dx)}{\pi_{\gamma^{*}}(B_{y})}c(x,u).$ (20)
Observe that the above is, see e.g. [15, Theorem 2.1], equal to the almost sure limit of the empirical expression on the right hand side below:
$C^{*}(y,u)=\lim_{N\to\infty}\frac{\sum_{k=0}^{N-1}c(X_{k},U_{k})1_{\{X_{k}\in B_{y},U_{k}=u\}}}{\sum_{k=0}^{N-1}1_{\{X_{k}\in B_{y},U_{k}=u\}}}$ (21)
Similarly, the effective transition probability $P^{*}(y^{\prime}|y,u)$ represents the probability of transitioning from bin $B_{y}$ to bin $B_{y^{\prime}}$ under action $u$, averaged over the invariant distribution:
$P^{*}(y^{\prime}|y,u)=\mathbb{P}_{x\sim\pi_{\gamma^{*}}|x\in B_{y}}[q(X_{t+1})=y^{\prime}|X_{t}=x,U_{t}=u]=\int_{B_{y}}\frac{\pi_{\gamma^{*}}(dx)}{\pi_{\gamma^{*}}(B_{y})}{\cal T}(B_{y^{\prime}}|x,u).$ (22)
Likewise, by [15, Theorem 2.1], the above is the almost sure empirical limit of of the right hand side below:
$P^{*}(y^{\prime}|y,u)=\lim_{N\to\infty}\frac{\sum_{k=0}^{N-1}1_{\{X_{k+1}\in B_{y^{\prime}}\}}1_{\{X_{k}\in B_{y},U_{k}=u\}}}{\sum_{k=0}^{N-1}1_{\{X_{k}\in B_{y},U_{k}=u\}}}$ (23)
An interpretation of the above result then is that one can first obtain the approximate model given with (20-22) by forcing the data into a Markovian model for both the empirical cost estimate (21) and empirical transition kernel estimate (23), and then solve the MDP as if this Markovian empirically constructed model is the actual one, instead of running Q-learning whose limit is then optimal precisely for this learned/empirically constructed model. A benefit of such a model-based approach is that one can have better sample complexity bounds compared with Q-learning for certain applications, see e.g. [35] (see also [13, Section 5.1]).
Remark 3.1.
In the planning framework presented in the previous section, we had the flexibility to select the weighting measures $\pi_{y_{i}}$ over the quantization bins arbitrarily. This allowed us to minimize the expected loss by choosing $\pi_{y_{i}}$ as a Dirac measure centered at the median of each bin. In contrast, within the learning context, the weighting measure $\pi_{y_{i}}$ is dictated by the exploration policy and is inherently dependent on the structure of the quantization bins $\{B_{i}\}$. Specifically, $\pi_{y_{i}}$ corresponds to the normalized invariant distribution over bin $B_{i}$ induced by the exploration policy; that is, it is the conditional probability measure given that the state is in $B_{i}$:
$\pi_{y_{i}}(A)=\frac{1_{\{A\in B_{i}\}}\pi_{\gamma^{*}}(A)}{\pi_{\gamma^{*}}(B_{y})},\qquad A\in B_{i},A\in{\cal B}(\mathbb{X})$
3.3 Error Analysis and Quantizer Design for Model Learning and Quantized Q-Learning
As noted above, the weighting measure depends on the exploration policy, quantizer and system model, and cannot be assigned arbitrarily. This dependence introduces a key difficulty in bounding the expected loss, as the weighing measure over the overflow bin, i.e. $\pi_{y_{M+1}}$ can no longer be freely chosen or controlled. This limitation is significant for the analysis of non-compact spaces. Nevertheless, we will show that it is still possible to establish a meaningful upper bound on the expected error by employing a Lyapunov-based analysis, analogous to the approach presented in Section 2.2.
We first start with the following result which relates the approximation error to occupation measures.
In the previous section, we studied the approximate model where the weighting measures $\hat{\pi}^{*}_{y_{i}}$ for each quantization bin were chosen freely and $\hat{\pi}^{*}_{y_{i}}$ were chosen as Dirac measures concentrated at the $\ell_{1}$ centroids (medians) of each bin $B_{i}$. However, when implementing Q-learning in the POMDP framework as described in this section, we lose the freedom to choose these weighting measures independently. The measures $\hat{\pi}^{*}_{y_{i}}$’s are inherently determined by the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy $\gamma^{*}$. We note that under weak continuity and compactness of the state space conditions, the set of invariant measures is compact in the weak topology. By considering this fact, we introduce the following theorem:
Theorem 13.
Under Assumption Assumption 2.2 and Assumption 3.1, given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}\in\mathcal{X}$:
$\left|J^{*}_{\beta}(x_{0})-\min_{v}\{Q^{*}(x_{0},v)\}\right|$
$\quad\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{B_{i}}\|x-x^{\prime}\|\mu_{\beta}(dx)\hat{\pi}_{y_{i}}(dx^{\prime}),$ (24)
where, $\Gamma_{s}$ represents the set of stationary policies, $\mu_{\beta}$ is the discounted occupation measure under the policy $\gamma_{s}$ and $\hat{\pi}_{y_{i}}(dx^{\prime})$ is the normalized invariant measure obtained under the exploration policy $\gamma^{*}$.
Proof.
From Theorem 2, we have the error bound:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\mathbb{E}_{x_{0}}^{\gamma_{s}}\left[\sum_{t=0}^{\infty}\beta^{t}L(X_{t})\right],$ (25)
where $\Gamma_{s}$ represents the stationary policies, and $L(X_{t})$ is defined in (5). As in the analysis leading to (B), substituting the expected accumulated discounted loss function, the bound is refined to (13), where $\hat{\pi}_{y_{i}}(dx^{\prime})$ is not freely chosen but obtained from the invariant probability measure under the exploration policy. ∎
In the following, we obtain an explicit bound which demonstrates the applicability of quantized Q-learning for non-compact spaces for quantized Q-learning.
Theorem 14.
Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
$\sup_{x\in\mathds{X},u\in\mathds{U}}\mathbb{E}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-\alpha V(x)+b,$ (26)
where $V(x)=|x|^{m}$, $m>1$. Then, under Assumption 2.1 and Assumption 2.2, provided that the cost function $c$ is bounded, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathbb{X}$:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (27)
where $C$ is a constant, which depends on the initial state $x_{0}$, defined as:
$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
Proof.
The proof is in Appendix G. ∎
4 Concluding Remarks
For MDPs with unbounded state spaces we presented upper bounds on finite model approximation errors via optimizing the quantizers used for finite model approximations. We also considered implications on quantized Q-learning and the performance of policies obtained via Q-learning where the quantized state is treated as the state itself. We noted the distinctions between planning, where approximating MDPs can be independently designed, and learning, where approximating MDPs are restricted to be defined by invariant measures of Markov chains under exploration policies, leading to significant subtleties on quantizer design performance. Nonetheless, asymptotic near optimality can be established under both setups with explicit convergence rates. In particular, under Lyapunov growth conditions, we obtained explicit upper bounds which decay to zero as the number of bins approaches infinity.
We note that, due to relative clarity in presentation especially involving the associated Lyapunov analysis, while we studied the case with the state space being $\mathbb{R}^{n}$, the analysis can be directly generalized to any normed space by adopting the required regularity conditions on the kernels and cost. Notably, if one applies the analysis here to a filter-reduced MDP (known as belief-MDP) of Partially Observable Markov Decision Processes (POMDPs) (see [12] for conditions on the necessary continuity properties), by replacing $\|\cdot\|_{1}$ with the Wasserstein distance of order-1 on probability measures, the analysis can be applied identically.
Appendix A Proof of Theorem 2
Proof.
We begin with the following initial bound, as in [16], using the corresponding Bellman equations:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)L(x_{0})+$
$\quad\beta\sup_{u_{0}\in\mathbb{U}}\mathbb{E}^{\gamma}\left[\left|J^{*}_{\beta}(X_{1})-\hat{J}_{\beta}(X_{1})\right|\big|x_{0},u_{0}\right],$ (28)
where $\alpha_{c}$ is the Lipschitz constant for the cost function, $\alpha_{T}$ is the Lipschitz constant for the transition kernel, and $L(x_{0})$ is the quantization error at $x_{0}$.
Let $V(x)=|J^{*}_{\beta}(x)-\hat{J}_{\beta}(x)|$ represent the difference between the optimal value function and the quantized value function. Then we have the following bound for $V(x_{0})$:
$V(x_{0})\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)L(x_{0})+\beta\sup_{u_{0}\in\mathbb{U}}\mathbb{E}\left[V(X_{1})\mid x_{0},u_{0}\right].$
One can show that the term $\mathbb{E}\left[V(X_{1})\mid x_{0},u_{0}\right]$ when considered as a function of $x_{0},u_{0}$ satisfies the measurable selection conditions under Assumption 2.2. We denote by $f(x)$ the control function which achieves the supremum such that
$\sup_{u_{0}}\mathbb{E}\left[V(X_{1})\mid x_{0},u_{0}\right]=\mathbb{E}\left[V(X_{1})\mid x_{0},f(x_{0})\right].$
Iterating the initial inequality for subsequent time step, we can write
$V(x_{0})$ $\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)\left(L(x_{0})+\beta\mathbb{E}_{x_{0}}\left[L(X_{1})|x_{0},f(x_{0})\right]\right)$
$\quad+\beta^{2}\mathbb{E}\left[\sup_{u_{1}}\mathbb{E}\left[V(X_{2})|X_{1},u_{1}\right]|x_{0},f(x_{0})\right]$
note that the supremum is achieved by the same control function $f$. Hence, defining the stationary policy $\gamma_{s}=\{f,f,f,\dots\}$, and repeating this process up to time step T1T-1, we have:
$V(x_{0})\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\sum_{t=0}^{T-1}\beta^{t}L(X_{t})\right]+\beta^{T}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[V(X_{T})\right].$
Since the cost function $c$ is bounded, it follows that $V(x)$ is bounded as well. Thus, $\lim_{T\to\infty}\beta^{T}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[V(X_{T})\right]=0$, which means the second term vanishes as $T\to\infty$. Taking the limit as $T\to\infty$, we get:
$V(x_{0})\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\sum_{t=0}^{\infty}\beta^{t}L(X_{t})\right].$
Finally, by taking the supremum over all stationary policies $\gamma_{s}$, we obtain the upper bound:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\beta\|\hat{J}_{\beta}\|_{\infty}\alpha_{T}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\sum_{t=0}^{\infty}\beta^{t}L(X_{t})\right].$
The proof is completed by noting that the term $\|\hat{J}_{\beta}\|_{\infty}$ is bounded by $\frac{\|c\|_{\infty}}{1-\beta}$, due to the boundedness of the cost function $c$. ∎
Appendix B Proof of Theorem 3
Proof.
The expected loss function $\mathbb{E}^{\gamma_{s}}[L(X_{t})]$ can be written as:
$\mathbb{E}^{\gamma_{s}}[L(X_{t})]=\sum_{i=1}^{M}\int_{B_{i}}L(x)\mu_{t}(dx),$
where $\mu_{t}(dx)$ is the distribution of the state $X_{t}$ at time $t$ which can also be seen as the marginal of the strategic measure $\mathds{P}^{\gamma_{s}}_{x_{0}}$ over $X_{t}$, and the summation is over all quantization bins $B_{i}$.
Next, we substitute the definition of $L(x)$ into the expectation:
$\mathbb{E}[L(X_{t})]=\sum_{i=1}^{M}\int_{B_{i}}\left(\int_{B_{i}}\|x-x^{\prime}\|_{1}\hat{\pi}_{y_{i}}(dx^{\prime})\right)\mu_{t}(dx).$
Next, we interchange the order of the integrals, which is justified due to the non-negativity of the terms by the Fubini–Tonelli theorem:
$\mathbb{E}[L(X_{t})]=\sum_{i=1}^{M}\int_{B_{i}}\int_{\mathds{X}}\mathds{1}_{B_{i}}(x)\|x-x^{\prime}\|_{1}\mu_{t}(dx)\hat{\pi}_{y_{i}}(dx^{\prime})$
where $\mathds{1}_{B_{i}}(x)$ is the indicator function.
Now, we return to the full expression for the error bound in (6). Then, $\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}[L(X_{t})]$ becomes:
$\sum_{i=1}^{M}\int_{B_{i}}\left(\sum_{t=0}^{\infty}\beta^{t}\int_{\mathds{X}}\mathds{1}_{B_{i}}(x)\|x-x^{\prime}\|_{1}\mu_{t}^{\gamma_{s}}(dx)\right)\hat{\pi}_{y_{i}}(dx^{\prime}).$ (29)
For a fixed $x^{\prime}$, define $\tilde{c}(x):=\mathds{1}_{B_{i}}(x)\|x-x^{\prime}\|_{1}$:
$\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\tilde{c}(X_{t})\right]$ $=\int_{\mathds{X}}\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\tilde{c}(X_{t})\mathds{1}_{\{X_{t}\in dx\}}\right]=\int_{\mathds{X}}\tilde{c}(x)\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\mathds{1}_{\{X_{t}\in dx\}}\right]$
$=\int_{\mathds{X}}\tilde{c}(x)\sum_{t=0}^{\infty}\beta^{t}\mathds{P}^{\gamma_{s}}_{x_{0}}\left(X_{t}\in dx\ \right)=\int_{\mathds{X}}\tilde{c}(x)\nu^{\gamma_{s}}_{x_{0}}(dx)=\frac{1}{1-\beta}\int_{\mathds{X}}\tilde{c}(x)\mu_{\beta}^{\gamma_{s}}(dx),$
where $\mu_{\beta}^{\gamma_{s}}(A):=(1-{\beta})\nu^{\gamma_{s}}_{x_{0}}(A)$ for $A\subseteq\mathds{X}$ and $\nu^{\gamma_{s}}_{x_{0}}$ is the discounted occupation measure as we defined earlier. We recognize this expression as a dot product between the cost function $\tilde{c}(x)$ and the normalized occupation measure $\mu_{\beta}^{\gamma_{s}}(dx)$, that is:
$\langle\mu_{\beta}^{\gamma_{s}},\tilde{c}\rangle=\int_{\mathds{X}}\tilde{c}(x)\mu_{\beta}^{\gamma_{s}}(dx),$
which leads to a linear program. Thus, the discounted sum can be expressed as:
$\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}^{\gamma_{s}}_{x_{0}}\left[\tilde{c}(X_{t})\right]=\frac{1}{1-\beta}\langle\mu_{\beta}^{\gamma_{s}},\tilde{c}\rangle$.
The full expression then, by considering the distribution on the realizations for $x^{\prime}$, for the discounted sum of the expected loss function becomes:
$\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{\mathds{X}}\|x-x^{\prime}\|_{1}\mathds{1}_{B_{i}}(x)\mu_{\beta}^{\gamma_{s}}(dx)\hat{\pi}_{y_{i}}(dx^{\prime})$
$=\frac{1}{1-\beta}\int_{\mathds{X}}\sum_{i=1}^{M}\mathds{1}_{B_{i}}(x)\int_{B_{i}}\|x-x^{\prime}\|_{1}\hat{\pi}_{y_{i}}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)=\frac{1}{1-\beta}\int_{\mathds{X}}L(x)\mu_{\beta}^{\gamma_{s}}(dx).$ (30)
Appendix C Proof of Theorem 7
Proof.
Step 1: Let $\mathds{X}\subseteq\mathbb{R}$ be the state space, which we partition into $M+1$ quantization bins $\{B_{1},B_{2},\ldots,B_{M},B_{M+1}\}$. The first $M$ bins cover a compact subset $\mathcal{K}=[-\frac{N}{2},\frac{N}{2}]\subset\mathds{X}$, and the last bin $B_{M+1}$ is an overflow bin that captures the rest of the state space outside $\mathcal{K}$.
We apply a uniform quantizer to the compact region $\mathcal{K}$, dividing it into $M$ bins of equal length. The quantization width of each bin is:
$\Delta=\frac{N}{M}.$
Step 2: For any state $x\in\mathcal{K}$, the quantization error $L(x)$ satisfies $L(x)\leq{\Delta}.$ We decompose the expected loss into two parts:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx),$ (31)
where $\mu_{\beta}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$.
Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
$\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}^{\gamma_{s}}(\mathcal{K})={\Delta}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (32)
Step 3: In the overflow bin $B_{M+1}$, the state space may be unbounded. In what follows, to simplify the analysis, we will assume that the overflow bin is always mapped to state $x=0$. We need to bound:
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}|x-x^{\prime}|\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(|x|+|x^{\prime}|\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)],$ (33)
where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized weighing measure over the bin $B_{M+1}$ and where the last step follows since we map the overflow bin directly to 0, i.e. $x^{\prime}=1$ with probability 1.
Using Hölder’s inequality to bound the expected loss over $B_{M+1}$:
$\int_{B_{M+1}}|x|\,\mu_{\beta}^{\gamma_{s}}(dx)=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]\leq\left(\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|^{m}]\right)^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (34)
Step 4: In the following, we bound the moment under the measure $\mu_{\beta}^{\gamma_{s}}$. Define the process $\{M_{t}\}$ for $t\geq 0$:
$M_{t}:=\frac{V(x_{t})}{(1-\alpha)^{t}}-\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}},$
with $M_{0}=V(x_{0})-b=|x_{0}|^{m}-b.$
Observe the following with respect to the filtration $\{\mathcal{F}_{t}\}$, where $\mathcal{F}_{t}=\sigma(X_{1},X_{2},\dots,X_{t})$ is the natural filtration:
$\mathbb{E}[M_{t+1}\mid\mathcal{F}_{t}]-M_{t}$ $=\mathbb{E}\left[\frac{V(X_{t+1})}{(1-\alpha)^{t+1}}-\sum_{k=0}^{t+1}\frac{b}{(1-\alpha)^{k}}\mid\mathcal{F}_{t}\right]-\frac{V(x_{t})}{(1-\alpha)^{t}}+\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
$=\frac{\mathbb{E}\left[V(X_{t+1})\mid\mathcal{F}_{t}\right]}{(1-\alpha)^{t+1}}-\sum_{k=0}^{t+1}\frac{b}{(1-\alpha)^{k}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
$\leq\frac{(1-\alpha)V(x_{t})}{(1-\alpha)^{t+1}}-\frac{b}{(1-\alpha)^{t+1}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\frac{b}{(1-\alpha)^{t+1}}$
$=0,$
where the inequality comes from the Lyapunov condition. Hence, we showed that the process $\{M_{t}\}$ is a supermartingale for all $t\geq 0$. Then, observe that for every fixed stopping time $t$:
$\mathbb{E}[M_{t}\mid\mathcal{F}_{0}]=\frac{\mathbb{E}\left[V(X_{t})\right]}{(1-\alpha)^{t}}-\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}\leq M_{0}=V(x_{0})-b.$
Rearranging the terms, we obtain:
$\mathbb{E}[V(X_{t})]$ $\leq(V(X_{0})-b)(1-\alpha)^{t}+(1-\alpha)^{t}\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
$=(V(X_{0})-b)(1-\alpha)^{t}+b\sum_{k=0}^{t}(1-\alpha)^{k}$
$=(V(X_{0})-b)(1-\alpha)^{t}+b\cdot\frac{1-(1-\alpha)^{t+1}}{\alpha}.$
Note that
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|^{m}]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[V(X)]$ $=(1-\beta)\sum_{t=0}^{\infty}\beta^{t}\mathbb{E}[V(X_{t})]$
$\leq(1-\beta)\sum_{t=0}^{\infty}\beta^{t}\left[(V(x_{0})-b)(1-\alpha)^{t}+\frac{b}{\alpha}(1-(1-\alpha)^{t+1})\right]$
$=(1-\beta)(V(x_{0})-b)\sum_{t=0}^{\infty}\left[\beta(1-\alpha)\right]^{t}+\frac{b(1-\beta)}{\alpha}\sum_{t=0}^{\infty}\beta^{t}\left[1-(1-\alpha)^{t+1}\right]$
$=\frac{(1-\beta)(V(x_{0})-b)}{1-\beta(1-\alpha)}+\frac{b(1-\beta)}{\alpha}\left[\frac{1}{1-\beta}-\frac{1-\alpha}{1-\beta(1-\alpha)}\right]$
$=\frac{V(x_{0})(1-\beta)+b\beta}{1-\beta(1-\alpha)}=:C,$ (35)
which shows that the error bound $C$ depends on the initial state $x_{0}$ and other constants.
Step 5: Returning back to equation (34), we find:
$\int_{B_{M+1}}|x|\,\mu_{\beta}^{\gamma_{s}}(dx)=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]$ $\leq\left(\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|^{m}]\right)^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}$
$\leq C^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$ (36)
where $C$ is defined in (35).
Combining the bounds from (32), (C) and (36), the total expected loss is bounded by:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq{\Delta}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+C^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (37)
Step 6: In particular, if $N$ is chosen to be $N=2(CM)^{\frac{1}{m}}$, then using Markov’s inequality, we obtain:
$\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mathbb{P}(|X|\geq N/2)\leq\frac{\mathbb{E}[|X|^{m}]}{\left((CM)^{1/m}\right)^{m}}\leq\frac{C}{(CM)}=\frac{1}{M}.$ (38)
and
$\Delta=\frac{N}{M}=\frac{2(CM)^{1/m}}{M}=\frac{2C^{1/m}}{M^{1-1/m}}.$
Using the bounds we obtained for the particular choice of $N$ in (37), we get:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq\frac{2C^{1/m}}{M^{1-1/m}}+C^{1/m}\cdot\left(\frac{1}{M}\right)^{1-1/m}\leq\frac{3C^{1/m}}{M^{1-1/m}}.$ (39)
With $m>1$, 11/m>01-1/m>0, and thus as $M\to\infty$, the total expected loss $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$. By combining this bound with Theorem 3, we obtain:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{3C^{1/m}}{(M^{1-1/m})(1-\beta)},$
where $M$ is the number of quantization bins, $m>1$ is the moment in the Lyapunov function, and $C$ is (defined in (35)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$.
Appendix D Proof of Theorem 8
Proof.
Step 1: Let the state space be $\mathds{X}\subseteq\mathbb{R}^{n}$ and fix a side–length parameter $N>0$. Choose an integer $k\geq 1$ and set
$M:=k^{n}\qquad\bigl(\text{so that }k=M^{1/n}\bigr).$
Define the centered nn–dimensional hyper-cube
$\mathcal{K}:=\Bigl[-\frac{N}{2},\frac{N}{2}\Bigr]^{n}\subset\mathds{X}.$
Partitioning each dimension of $\mathcal{K}$ with k bins uniformly yields a quantization bin with width:
$\Delta=\frac{N}{k}=\frac{N}{M^{1/n}},$
thus producing $M$ congruent hyper-cubic cells of volume $\Delta^{\,n}=N^{n}/M$. Index these cells by
$B_{i_{1},\dots,i_{n}}\;=\;\prod_{j=1}^{n}\Bigl[-\tfrac{N}{2}+(i_{j}-1)\Delta,\;-\tfrac{N}{2}+i_{j}\Delta\Bigr),\quad i_{j}\in\{1,\dots,k\},$
and enumerate them as $\{B_{1},\dots,B_{M}\}$. Set
$B_{M+1}\;:=\mathds{X}\setminus\mathcal{K},$
which is the set of states outside the compact (granular) grid.
Step 2: For any state $x\in\mathcal{K}$, due to the uniform quantization, we have that
$L(x)\leq{n\Delta}$
We decompose the expected loss into two parts as in the proof of Theorem 7:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\mu_{\beta}^{\gamma_{s}}(dx),$ (40)
where $\mu_{\beta}^{\gamma_{s}}$ is the normalized discounted occupation measure under the stationary policy $\gamma_{s}$.
Inside $\mathcal{K}$ the error is uniformly bounded, hence
$\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)\leq{n\,\Delta\,\mu_{\beta}^{\gamma_{s}}\!(\mathcal{K})}={n\,\Delta\bigl[1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\bigr]}.$ (41)
Step 3 : Inside the overflow bin $B_{M+1}\subseteq\mathds{X}$ the state space is unbounded. In the quantizer, for simplicity in analysis, we may take that the overflow bin is directly mapped to state $x=0$. We must bound:
$\int_{B_{M+1}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}\|x-x^{\prime}\|_{1}\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\bigl(\|x\|_{1}+\|x^{\prime}\|_{1}\bigr)\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]$
where the last step follows since we assume that $x^{\prime}=0$ always for the overflow bin. We apply Hölder’s inequality with exponent $m>1$:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]\leq\Bigl(\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}^{m}\bigr]\Bigr)^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (42)
Step 4: Repeating the arguments in Step 4 of the proof of Theorem 7, we can show that
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]\leq\frac{V(x_{0})(1-\beta)+b\beta}{1-\beta(1-\alpha)}=:C$
Combining (41) with the overflow estimate obtained in (42), we get:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq{n\Delta\bigl[1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\bigr]}+C^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (43)
We select $N\;=\;2\bigl(C\,M\bigr)^{1/m}.$ By Markov’s Inequality:
$\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $={\mu_{\beta}^{\gamma_{s}}(\{x={x_{1},\cdots,x_{n}}:\min_{i=1,\cdots,n}|x_{i}|\geq N/2\})\leq}\mathop{\rm Pr}\bigl[\|X\|_{1}\geq N/2\bigr]$
$\leq\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]}{(\tfrac{N}{2})^{m}}=\frac{C}{C\,M}=\frac{1}{M}.$
With $k=M^{1/n}$ interior cells per axis, the hyper-cube side length is
$\Delta=\frac{N}{k}=2C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}.$
Inserting these bounds into (43) and assuming that $m>n$:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}\left(\frac{1}{M}\right)^{1-\frac{1}{m}}$
$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{-\left(1-\frac{1}{m}\right)}$
$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
$=(2n+1)C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
$=(2n+1)C^{1/m}M^{-\left(\frac{1}{n}-\frac{1}{m}\right)}.$
As a result, as $M\to\infty$, the total expected loss $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$ since $m>n$.
Under necessary conditions, we obtain:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{1/n-1/m})(1-\beta)},$ (44)
where $n$ is the dimension of the state space, $M$ is the number of quantization bins, $m>n$ is the moment in the Lyapunov function, and $C$ is (defined in (35)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$. ∎
Appendix E Proof of Theorem 9
Proof.
Using Assumption 2.3, the following Average Cost Optimality Equation (ACOE) is satisfied for the original model:
$h(x)=\inf_{u\in\mathds{U}}\left\{c(x,u)+\int_{\mathds{X}}h(x_{1})\mathcal{T}(dx_{1}|x,u)\right\}-\int_{\mathds{X}}h(x)\mu(dx).$ (45)
Similarly, for the discretized model, we have the corresponding ACOE:
$\hat{h}(y)=\inf_{u\in\mathds{U}}\left\{C^{*}(y,u)+\sum_{y_{1}}\hat{h}(y_{1})P^{*}(y_{1}|y,u)\right\}-\int_{\mathds{X}}\hat{h}(q(x))\mu(dx).$ (46)
For the rest of the proof, we simply use the same notation for $C^{*}$, and $\hat{h}$ by extending them as constant over the quantization bins, that is we override the notation and use $\hat{h}(x)$ for $\hat{h}(q(x))$ and $C^{*}(x,u)$ for $C^{*}(q(x),u)$.
By defining a version of the finite model kernel $P^{*}$ that is defined over $\mathds{X}$ such that
$\hat{\mathcal{T}}(\cdot|x,u):=\int_{B_{i}}\mathcal{T}(\cdot|x^{\prime},u)\hat{\pi}_{i}(dx^{\prime})\text{ for }x\in B_{i},$
one can show that
$\hat{h}(x)=\inf_{u\in\mathds{U}}\left\{C^{*}(x,u)+\int\hat{h}(x_{1})\hat{\mathcal{T}}(dx_{1}|x,u)\right\}-\int_{\mathds{X}}\hat{h}(x)\mu(dx).$ (47)
According to Verification Theorem for average cost criterion (see [9] and [1]), the long-term average cost $J^{*}(x)$ for the original model and $\hat{J}(x)$ for the discretized model are given by:
$J^{*}(x)=j^{*}=\int_{\mathds{X}}h(x)\mu(dx),$
$\hat{J}(x)=\hat{j}=\int_{\mathds{X}}\hat{h}(x)\mu(dx),$
[…]
We denote by $\mathcal{T}^{-}(\cdot|x,u):=\mathcal{T}(\cdot|x,u)-\mu(\cdot)$ and $\hat{\mathcal{T}}^{-}(\cdot|x,u):=\hat{\mathcal{T}}(\cdot|x,u)-\mu(\cdot)$. Note that
$\|\mathcal{T}^{-}(\cdot|x,u)-\hat{\mathcal{T}}^{-}(\cdot|x,u)\|_{TV}=\|\mathcal{T}(\cdot|x,u)-\hat{\mathcal{T}}(\cdot|x,u)\|_{TV}.$
We further denote by $V(x):=|h(x)-\hat{h}(x)|$. In what follows, the term $\sup_{u}\int V(x_{1})\mathcal{T}(dx_{1}|x,u)$ will be of interest. We will denote the control function that achieves the supremum by $\gamma_{s}$, whose existence is guaranteed under Assumption 2.2. Comparing the ACOEs for corresponding models, we can write that
$V(x)$ $\leq\sup_{u\in\mathds{U}}\bigg|(c(x,u)-C^{*}(x,u))+\left(\int h(x_{1})\mathcal{T}^{-}(dx_{1}|x,u)-\int\hat{h}(x_{1})\hat{\mathcal{T}}^{-}(dx_{1}|x,u)\right)\bigg|$
$\leq\alpha_{c}L(x)+\sup_{u\in\mathds{U}}\bigg|\int h(x_{1})\mathcal{T}^{-}(dx_{1}|x,u)-\int\hat{h}(x_{1})\mathcal{T}^{-}(dx_{1}|x,u)\bigg|$
$\qquad\qquad+\sup_{u\in\mathds{U}}\bigg|\int\hat{h}(x_{1})\mathcal{T}^{-}(dx_{1}|x,u)-\int\hat{h}(x_{1})\hat{\mathcal{T}}^{-}(dx_{1}|x,u)\bigg|$
$\leq\alpha_{c}L(x)+\sup_{u\in\mathds{U}}\int V(x_{1})\mathcal{T}^{-}(dx_{1}|x,u)+\|\hat{h}\|_{\infty}\alpha_{T}L(x)$
$=(\alpha_{c}+\|h\|_{\infty}\alpha_{T})L(x)+\int V(x_{1})\mathcal{T}(dx_{1}|x,\gamma_{s}(x))-\int V(x)\mu(dx).$
By repeating the same step, one can write that for any $T<\infty$:
$V(x)\leq(\alpha_{c}+\|h\|_{\infty}\alpha_{T})\sum_{t=0}^{T-1}E^{\gamma_{s}}\left[L(X_{t})\right]-T\int V(x)\mu(dx).$
We can then write that
$\int V(x)\mu(dx)+\frac{V(x)}{T}\leq(\alpha_{c}+\|h\|_{\infty}\alpha_{T})\frac{1}{T}\sum_{t=0}^{T-1}E^{\gamma_{s}}\left[L(X_{t})\right].$
At the end of the proof, we will show that $h(x),\hat{h}(x)$ and thus $V(x)$ are uniformly bounded. Assuming this is true and sending $T\to\infty$, we get
$\int V(x)\mu(dx)+\frac{V(x)}{T}$ $\leq(\alpha_{c}+\|h\|_{\infty}\alpha_{T})\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}E^{\gamma_{s}}\left[L(X_{t})\right]$
$=(\alpha_{c}+\|h\|_{\infty}\alpha_{T})\int L(x)\pi_{\gamma_{s}}(dx),$
where $\pi_{\gamma_{s}}$ is the invariant measure induced by policy $\gamma_{s}$ which exists under Assumption 2.3.
To show the boundedness, we let $\hat{T}$ denote the ACOE operator for the finite model. Starting from $h_{0}(x)=0$, let $h_{k}$ denote the function we obtain after applying the operator $\hat{T}$, $k$ consecutive times. We then have that
$h_{k}(x)\leq\|c\|_{\infty}+\sup_{u}\left\{\int h_{k-1}(x_{1})\hat{\mathcal{T}}^{-}(dx_{1}|x,u)\right\}\leq\|c\|_{\infty}+\|h_{k-1}\|_{\infty}(1-\mu(\mathds{X})).$
Letting $\alpha:=(1-\mu(\mathds{X}))$, and repeating this step, we write
$\|h_{k}\|_{\infty}\leq\|c\|_{\infty}\sum_{t=0}^{k-1}\alpha^{t}.$
Finally, using the fact that the operator $\hat{T}$ is a contraction under the supremum norm under Assumption 2.3 with the fixed point $\hat{h}$, we conclude
$\|\hat{h}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\alpha}=\frac{\|c\|_{\infty}}{\mu(\mathds{X})}.$
Identical steps can be used to show that $\|h\|_{\infty}\leq\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$. Combining what we have so far, we write
[…]
$\leq\left(\alpha_{c}+\frac{\|c\|_{\infty}\alpha_{T}}{\mu(\mathds{X})}\right)\int L(x)\pi_{\gamma_{s}}(dx)$
which completes the proof.
Appendix F Proof of Theorem 11
Proof.
First we assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
$\sup_{x\in\mathds{X},u\in\mathds{U}}\mathbb{E}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$ (48)
where $f(x)=|x|^{m}$, $m>1$.
Then, we choose a collection of quantization bin $\{B_{i}\}_{i}^{M+1}$ such that the first $M$ bins quantize the compact set $(bM)^{1/m},(bM)^{1/m}]$ uniformly and the last bin $B_{M+1}$ is the overflow bin that captures the rest of the state space. Then, the total expected distortion becomes:
$\mathbb{E}_{\pi_{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\,\pi_{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\,\pi_{\gamma_{s}}(dx)$
In the overflow bin $B_{M+1}$, the state space may be unbounded, and the quantization error $L(x)=\|x-y_{M+1}\|$ can be large. We need to bound:
$\int_{B_{M+1}}L(x)\,\pi_{\gamma_{s}}(dx)=\int_{B_{M+1}}|x-y_{M+1}|\,\pi_{\gamma_{s}}(dx),$ (49)
where $y_{M}$ is the representative point for bin $B_{M+1}$.
We choose $y_{M+1}=0$ (or any fixed point), so that $L(x)=|x|$ for $x\in B_{M+1}$.
By our assumption, there exists a Lyapunov function $V$ that satisfies inequality (48). Then, by Theorem 10, for any initial state $x_{0}=x$, we have:
$\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq V(x)+\mathbb{E}\left[\sum_{t=0}^{T-1}b\right]=V(x)+bT,$
where $T$ is a deterministic stopping time. By taking the limit for both sides, we get:
$\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq\limsup_{T\to\infty}\frac{1}{T}(V(x)+bT)=b,$
Now, consider any invariant probability measure $\pi$. Recall that we assigned $f=|x|^{m},m>1$. Fix $N>0$ and define $f_{N}(x)=\min(|x|^{m},N)$. Then, we can write (see [34, p. 75]):
$\int_{\mathds{X}}f_{N}(x)\pi(dx)$ $=\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{x_{0}\sim\pi}\left[E_{x_{0}}[f_{N}(X_{t})]\right]$
$\leq\mathbb{E}_{\pi}\left[\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}E_{x_{0}}[f_{N}(X_{t})]\right]\leq b,$
where the first inequality comes from applying Fatou’s Lemma since $f_{N}(x)$ is bounded by $N$. By Taking the limit $N\to\infty$ and noting that $f_{N}(x)\to|x|^{m}$ pointwise and monotonically, by Monotone Convergence Theorem, we obtain $\int_{\mathds{X}}|x|^{m}\pi(dx)\leq b$.
Finally, since this holds for any invariant probability measure $\pi$ and $|x|^{m}$ is non-negative, we conclude that $\mathbb{E}[|X_{t}|^{m}]\leq b,\forall t\geq 0.$ By following the same procedure in the proof of Theorem 7, we obtain the following bound:
$\mathbb{E}_{\pi_{\gamma_{s}}}[L(X)]$ $\leq\Delta\cdot\left(1-\pi_{\gamma_{s}}(B_{M+1})\right)+b^{1/m}\cdot\pi_{\gamma_{s}}(B_{M+1})^{1-1/m}$ (50)
$\leq\Delta\ +\frac{b^{1/m}}{M^{1-1/m}}\leq\frac{2b^{1/m}}{M^{1-1/m}}+\frac{b^{1/m}}{M^{1-1/m}}=\frac{3b^{1/m}}{M^{1-1/m}}.$ (51)
By combining this bound with Theorem 9 under the given assumptions, we obtain:
$\left|\hat{J}_{avg}(x_{0})-J^{*}_{avg}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{(M^{1-1/m})},$ (52)
where $M$ is the number of quantization bins, $m>1$ is the moment in Lyapunov function and $b$ is the uniform bound on the $m$-th moment of $X_{t}$. Thus, the average cost leads to sharper bounds via the Foster-Lyapunov analysis. ∎
Appendix G Proof of Theorem 14
Proof.
Step1: We partition the state space similar to the quantization in Theorem 7 by choosing the compact set $\mathcal{K}=[-(CM)^{1/m},(CM)^{1/m}]$. Then, for any state $x\in\mathcal{K}$, the quantization error $L(x)$ satisfies:
$L(x)\leq{\Delta}=\frac{2(CM)^{1/m}}{M}=\frac{2C^{1/m}}{M^{1-1/m}}.$ (53)
We decompose the total expected loss into two parts:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx),$ (54)
where $\mu_{\beta}^{\gamma_{s}}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$.
Step 2: Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
$\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}(\mathcal{K})=\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (55)
Step 3 (Two parallel arguments): In the overflow bin $B_{M+1}$, the state space may be unbounded. Unlike the analysis in Section 2.2, we cannot assign arbitrary weighting measure $\hat{\pi}_{M+1}$ for the overflow bin. Thus, the bound becomes:
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}|x-x^{\prime}|\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (56)
$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(|x|+|x^{\prime}|\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (57)
$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]+\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1}),$ (58)
where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized invariant measure of the state process over the bin $B_{M+1}$ under the exploration policy. The second term is handled in two ways:
3.a
Observe that if $x_{0}\sim\pi_{\gamma^{*}}$ then $\mu_{\beta}^{\gamma_{s}}=\pi_{\gamma^{*}}$ and therefore the terms in the summation above, that is,
$\mathbb{E}_{\hat{\pi}^{*}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mu_{\beta}^{\gamma_{s}}(B_{M+1})\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]$
will be identical.
Similar to the proof of Theorem 7, by the law of the iterated expectations, we can show that when the initial state is not fixed but given by $x_{0}\sim\pi_{\gamma^{*}}$:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\mathbb{E}[|X|\mathds{1}_{B_{M+1}}(X)|X_{0}=x]]\leq\hat{C}^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$
where
$\hat{C}:=\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|x_{0}|]^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
By using Markov’s inequality, we obtain:
$\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mathbb{P}\left(|X|\geq(CM)^{1/m}\right)\leq\frac{\mathbb{E}[|X|^{m}]}{\left((CM)^{1/m}\right)^{m}}\leq\frac{1}{M},$ (59)
which gives:
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2\hat{C}^{1/m}}{M^{1-1/m}}$ (60)
3.b
We know that $\hat{\pi}_{M+1}$ is the normalized measure of the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy. Then, we have:
$\mathbb{E}_{\hat{\pi}_{M+1}}[|X|]=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}\cdot\pi_{\gamma^{*}}(B_{M+1})^{1-1/m}}{\pi_{\gamma^{*}}(B_{M+1})}=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}$
By using Comparison Theorem 10 and using a similar procedure in Section 2.3, under the Lyapunov condition in (26) we obtain:
$\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]\leq\frac{b}{\alpha}.$ (61)
For the overflow set $B_{M+1}=\{|x|\geq(CM)^{1/m}\}$ we have by Markov’s inequality:
$\pi_{\gamma^{*}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]}{(CM)^{m/m}}\leq\frac{b/\alpha}{CM}.$
Putting the pieces together:
$\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}\cdot\frac{1}{M}$
$=\Bigl(\tfrac{b}{\alpha}\Bigr)^{1/m}\Bigl(\tfrac{C}{b/\alpha}\Bigr)^{1/m}M^{-(1-1/m)}=\frac{C^{1/m}}{M^{1-1/m}}$
and
$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2C^{1/m}}{M^{1-1/m}},$
which goes to zero as M goes to infinity. Thus the expected quantization error for the overflow bin vanishes as the number of quantization bin, M, increases.
Either method gives the same final error bound stated in Theorem 14. While the first method assumes that the initial state distribution is the same as $\pi_{\gamma^{*}}$, the second method assumes that the initial state can be any deterministic state in $\mathbb{X}$.
Step 4: Now, combining the bounds we obtained so far, we get:
$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+\frac{2C^{1/m}}{M^{1-1/m}}$
$\leq\frac{2C^{1/m}}{M^{1-1/m}}+\frac{2C^{1/m}}{M^{1-1/m}}=\frac{4C^{1/m}}{M^{1-1/m}}.$ (62)
Obviously, $m>1$ implies 11/m>01-1/m>0.
Step 5: Let By combining this bound with Theorem 3 under the necessary assumptions, we obtain:
$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (63)
where $M$ is the number of quantization bins, $m>1$ is the moment in Lyapunov function, $C$ is the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}$. ∎

### v2

@@
 To this end, the goal of this paper is to design quantizers in such a finite model approximation and learning framework for continuous space MDPs.
-Notably, this paper develops a mathematical framework and formulation paving the way for non-uniform quantizer design and associated performance bounds, which are then also applicable to MDPs with non-compact spaces.
 1.1 Related Literature
@@
 However, learning optimal Q-functions with linear function approximation is known to be unstable in general [2] except in special cases.
-For general linear function approximation, [20] has shown that under certain a class of exploration policies, optimal Q-value learning with linear function approximation remains bounded.
+For general linear function approximation, [19] has shown that under a certain class of exploration policies, optimal Q-value learning with linear function approximation remains bounded.
 The special cases where the convergence can be guaranteed include when (i) the exploration policy is already close to the greedy policy of the learning iterations [17], and (ii) the stage-wise cost function, the transition model, and thus the optimal Q-functions are perfectly represented by the basis functions, i.e.
@@
 In particular, [15] showed that the Q-values learned under quantization based learning correspond to an auxiliary finite control problem, which is a finite approximation of the original control problem with a particular weight measure on the quantization bins.
-This observation has an important consequence that the model-based approaches using state space quantization and the quantized Q-learning coincide and can be used interchangeably, even under very mild continuity conditions on the transition kernels [27].
+This observation has an important consequence that the model-based approaches using state space quantization and the quantized Q-learning coincide and can be used interchangeably, even under mild continuity conditions on transition kernels [26].
 For quantization based learning methods while often convergence is studied, error analysis regarding the limit of the stochastic iterates is typically not studied in general or an error analysis is not provided at all.
@@
 Another direction of research with provable loss guarantees is kernel-based methods [21, 20, 3, 6, 34, 7] where the cost and the transition models are estimated via empirical kernel regression or otherwise simulations are used to update the Bellman optimality equations.
-One can then find the control policies based on the learned models using model-based approaches.
-Most prior work focuses on approximation of control problems with compact and bounded spaces.
+One can then find control policies based on the learned models using model-based approaches.
+Most prior work focuses on approximation of control problems with compact spaces.
 There is a limited number of provable guarantees for the approximation of control problems with unbounded state spaces.
@@
 1.3 Contributions
-In this paper, we make the following contributions:
-  • (i)
-    In Theorem 2, we refine and computationally improve the upper bounds given in [16] which involve admissible policies to ones that only involve stationary policies.
+In this paper, we make the following contributions: (i) In Theorem 2, we refine and computationally improve the upper bounds given in [15] which involve admissible policies to ones that only involve stationary policies.
 This facilitates an analysis involving occupation measures as well as invariant measures for discounted and ergodic cost criteria, respectively; which is then utilized later in the paper.
-  • (ii)
-    In Theorems 3 9, we derive bounds in terms of occupation measures, to represent the loss in terms of occupation and invariant measures, respectively for the discounted and average cost criteria.
+(ii) In Theorems 3 7, we derive bounds in terms of occupation measures, to represent the loss in terms of occupation and invariant measures, respectively for the discounted and average cost criteria.
 In Corollary 5, we optimize the quantization design by choosing the representative points as medians (in the $\ell_{1}$ sense) with respect to the occupation measures and arrive at an explicit error bound.
 We thus provide a convergence rate analysis of quantization based approximations which is convenient for stochastic analysis and which is also applicable for non-compact state spaces.
-  • (iii)
-    In Theorems 7 and 8 under Foster-Lyapunov conditions, we derive explicit error bounds which decay to zero as the quantization gets finer and show better performance than uniform quantization in MDPs with non-compact state spaces under discounted cost criteria.
+(iii) In Theorem 6, under Foster-Lyapunov conditions, we derive explicit error bounds which decay to zero as the quantization gets finer and show better performance than uniform quantization in MDPs with non-compact state spaces under discounted cost criteria.
 We extend our analysis to the average cost criterion by obtaining explicit error bounds in Theorem 8.
-  • (iv)
-    In Theorem 14, we extend our analysis on quantizer design to quantized Q-learning and empirical model learning (which are equivalent in performance).
+(iv) In Theorem 11, we extend our analysis on quantizer design to quantized Q-learning and empirical model learning (which are equivalent in performance).
 We obtain the error bounds which depend, unlike the planning problem above, on the invariant measure induced by the exploration policy in Theorem 10.
@@
 The above ensures that optimal policies exist.
-Furthermore, by [24, Lemma 3.19] and [24, Theorem 3.16] (see also [26]), any MDP with a weakly continuous transition probability kernel can be approximated by an MDP with finite action spaces.
+Furthermore, by [23, Lemma 3.19] and [23, Theorem 3.16] (see also [25]), any MDP with a weakly continuous transition probability kernel can be approximated by an MDP with finite action spaces.
 Accordingly, in the sequel, we assume that $\mathds{U}$ is finite.
-Remark 2.2.
-[…]
 Assumption 2.2.
@@
 $\nu_{x_{0}}^{\gamma_{s}}(D)=\sum_{t=0}^{\infty}\beta^{t}\,\mathbb{E}_{x_{0}}^{\gamma_{s}}\!\bigl[\mathbf{1}_{D}(X_{t},U_{t})\bigr]=\sum_{t=0}^{\infty}\beta^{t}\,\mathbb{P}_{x_{0}}^{\gamma_{s}}\!\bigl((X_{t},U_{t})\in D\bigr),$
-where the probability measure $\mathbb{P}_{x_{0}}^{\gamma_{s}}$ over the state and action process is defined by the initial condition $x_{0}$, the stationary policy $\gamma_{s}$, and the transition kernel $\mathcal{T}$.
+where the probability measure $\mathbb{P}_{x_{0}}^{\gamma_{s}}$ over the state and action process is defined by the initial condition $x_{0}$, the policy $\gamma_{s}$, and the kernel $\mathcal{T}$.
 For a product set $A\times\mathds{U}$ with $A\subseteq\mathds{X}$ measurable we obtain
@@
 The proof is in Appendix B. ∎
-To minimize the error bound obtained, we optimize the quantizer by choosing the centroids of the quantization bins properly. The following lemma will be useful in choosing the centroids.
+The following is a supporting result.
 Lemma 4.
-[5] Let $\mathds{X}\subseteq\mathds{R}$, and let $X$ be a random variable with distribution $\mu$ over a quantization bin $B_{i}$.
-Suppose that $\mathbb{E}(|X|)<\infty$.
-Then, the value $y_{i}$ that minimizes the expected distortion $\mathbb{E}[|X-y_{i}|]$ is the median of $X$ with respect to the probability measure $\mu$ over $B_{i}$ .
-For a proof, see [5].
-According to Lemma 4, we establish that for a given quantization bin $B_{i}$, the expected distortion is minimized when the representative point $y_{i}$ is chosen as the median of $B_{i}$ with respect to the probability measure $\mu$ over $B_{i}$.
-This turns the normalized measure $\hat{\pi}_{y_{i}}$ into the Dirac measure $\delta_{y_{i}}$ concentrated at $y_{i}$.
-This result can be also applied to higher dimensional state spaces in $\mathbb{R}^{n}$ according to the following theorem.
-The result can be extended to higher dimensions when considering the $\ell_{1}$ norm in $\mathbb{R}^{n}$.
-Lemma 5.
-Let $X=(X_{1},X_{2},\dots,X_{n})$ be a random vector with finite expected absolute deviations in each coordinate.
+Let $X=(X_{1},X_{2},\dots,X_{n})$ be a $\mathds{R}^{n}$-valued random vector with finite expected absolute deviations in each coordinate.
 Then, the point $y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*})$ that minimizes $\mathbb{E}[\|X-y_{i}\|_{1}]$ is obtained by choosing each $y_{i,k}^{*}$ to be a median of the marginal distribution of $X_{k}$ over the bin $B_{i}$, for $k=1,2,\dots,n$.
-The proof of the above is immediate: Let $X$ be an n -dimensional random vector, and $y_{i}\in\mathbb{R}^{n}$ . The expected $\ell_{1}$ distortion is:
+The case with $n=1$ is proven in [5]. The generalization for the $n$-dimensional case is then immediate: Let $y_{i}\in\mathbb{R}^{n}$ . The expected $\ell_{1}$ distortion writes as:
 $\mathbb{E}[\|X-y_{i}\|_{1}]=\int_{B_{i}}\|x-y_{i}\|_{1}\mu(dx)=\sum_{k=1}^{n}\int_{B_{i}}|x_{k}-y_{i,k}|\mu_{k}(dx),$ (7)
-where $\mu_{k}$ is the marginal distribution of $X_{k}$ over the bin $B_{i}$.
-Since the sum of convex functions is convex, $\mathbb{E}[\|X-y_{i}\|_{1}]$ is convex in $y_{i}$ .
-Each component $y_{i,k}$ appears only in the term involving $x_{k}$, so we can minimize each term separately.
-By the same reasoning as in the scalar case, as in Lemma 4, $\int_{B_{i}}|x_{k}-y_{i,k}|\mu_{k}(dx)$ is minimized when $y_{i,k}$ is a median of $X_{k}$ with respect to the probablity measure $\mu_{k}$ over $B_{i}$.
-Therefore, the point $y^{*}_{i}$ that minimizes the expected $\ell_{1}$ distortion is found as:
-$y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*}),\quad\text{for}\quad k=1,2,\dots,n$ (8)
-where each $y_{i,k}^{*}$ is the median of the quantization bin $B_{i}$ with respect to the marginal distribution $\mu_{k}$.
+where $\mu_{k}$ is the marginal distribution of $X_{k}$ over the bin $B_{i}$. We thus can minimize each term separately and the result follows.
 Corollary 5 (to Theorem 3 and Theorem 4).
@@
 Theorem 6.
-Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $\alpha>0$ and $V:\mathds{X}\to[0,\infty)$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
-$\mathbb{E}^{\gamma_{s}}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-\alpha V(x)+b,\quad x\in\mathds{X},u\in\mathds{U}$ (9)
-for all $\gamma_{s}\in\Gamma_{s}$ where $V(x)=|x|^{m}$, for some $m>1$.
-Then, under Assumption 2.1 and Assumption 2.2, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathds{X}$:
-$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{3C^{1/m}}{(M^{1-1/m})(1-\beta)},$
-where $C$ is a constant, which depends on the initial state $x_{0}$, defined as:
-$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
+Let $\mathds{X}\subseteq\mathbb{R}^{n}$, $b\geq 0$, $\alpha>0$, and define the Lyapunov function $V(x)=\|x\|_{1}^{m}.$ Assume the controlled process $\{X_{t}\}$ satisfies the drift condition
+$\mathbb{E}\bigl[V(X_{t+1})\mid X_{t}=x,U_{t}=u\bigr]\leq V(x)-\alpha V(x)+b,\qquad x\in\mathds{X},u\in\mathds{U}$
+Under Assumption 2.1 and Assumption 2.2, let $M$ be the total number of hyper-cubic bins in the uniform quantizer. Then, for every initial state $x_{0}\in\mathds{X}$ we have that,
+$\bigl|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\bigr|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{\frac{1}{n}(1-\frac{1}{m})})(1-\beta)},$ (8)
+where
+$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
 Proof.
@@
 Observe that as $M\to\infty$, the error converges to zero.
-The next result extends Theorem 7 to an nn–dimensional state space $\mathds{X}\subseteq\mathbb{R}^{n}$ (with the $\ell_{1}$–norm).
-Theorem 8.
-Let $b\geq 0$, $\alpha>0$, and define the Lyapunov function $V(x)=\|x\|_{1}^{m}$ with exponent $m>n$. Assume the controlled process $\{X_{t}\}$ satisfies the drift condition
-$\mathbb{E}^{\gamma_{s}}\bigl[V(X_{t+1})\mid X_{t}=x,U_{t}=u\bigr]\leq V(x)-\alpha V(x)+b,\qquad x\in\mathds{X},u\in\mathds{U}$
-for all $\gamma^{s}\in\Gamma_{s}$.
-Under Assumption 2.1 and Assumption 2.2, let $M$ be the total number of hyper-cubic bins in the uniform quantizer.
-Then, for every initial state $x_{0}\in\mathds{X}$ we have that,
-$\bigl|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\bigr|\leq\Bigl(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\Bigr)\frac{(2n+1)C^{1/m}}{M^{\frac{1}{n}-\frac{1}{m}}(1-\beta)},$ (10)
-where
-$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
-Proof.
-The proof is in Appendix D. ∎
-The explicit quantizer is presented in the proof of Theorem 8.
 2.3 Refined Error Bounds for the Average Cost Criterion
@@
 Proof.
-The proof is in Appendix E
-Now, we state the following theorem, which will be useful in the proof of our next result:
-Theorem 10 (Comparison Theorem).
-[19, Theorem 14.2.2] Let $V:\mathds{X}\to[0,\infty)$ and $f,g:\mathds{X}\to[0,\infty)$. Let $\{x_{n}\}$ be a Markov chain on $\mathds{X}$. If the following condition is satisfied:
-$\int_{\mathds{X}}P(x,dy)V(y)\leq V(x)-f(x)+g(x),\quad x\in\mathds{X},$
-then, for any stopping time $\tau$ with $\mathbb{P}(\tau<\infty)=1$, it holds that
-$\mathbb{E}\left[\sum_{t=0}^{\tau-1}f(x_{t})\right]\leq V(x_{0})+\mathbb{E}\left[\sum_{t=0}^{\tau-1}g(x_{t})\right].$
-Proof.
-See [34]. ∎
-Now, we present our result under Lyapunov condition similar to the quantizer design for discounted cost criterion.
+The proof is in Appendix D
 Theorem 8.
-Assume $\mathds{X}\subseteq\mathbb{R}$, $f(x)=\|x\|^{m}$ with $m>1$, and suppose
-$\sup_{x,u}\mathbb{E}[V(X_{t+1})\mid X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$
+Assume $\mathds{X}\subseteq\mathbb{R}^{n}$, $f(x)=\|x\|_{1}^{m}$ with $m>1$, and suppose
+$\sup_{x,u}\mathbb{E}[V(X_{t+1})\mid X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$ (11)
 where $b\geq 0$ and $V:\mathds{X}\to[0,\infty)$ is a Lyapunov function.
 Then, under Assumption 2.1, Assumption 2.2 and Assumption 2.3, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathds{X}$:
-$|\hat{J}_{avg}(x_{0})-J_{avg}^{*}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{M^{1-1/m}}.$ (13)
+$|\hat{J}_{avg}(x_{0})-J_{avg}^{*}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{(2n+1)b^{1/m}}{M^{1/n(1-1/m)}}.$ (12)
 Proof.
-The proof is in Appendix F
+The proof is in Appendix E
 As earlier, as $M\to\infty$ the error converges to zero.
@@
 As noted in [13], quantizing the state space of a continuous MDP converts the problem into a Partially Observable Markov Decision Process (POMDP) which is then non-Markovian.
-In particular, with the quantization process, the controller has partial observation.
-Define
-$Y_{t}=q(X_{t}),$
-where $q:\mathbb{X}\rightarrow\mathbb{Y}$ is the quantizer, which maps the continuous state space $\mathbb{X}$ to the finite set of representative points $\mathbb{Y}=\{y_{1},y_{2},\ldots,y_{M}\}$.
-The observation model in our setup is described as:
-$O(y_{i}\mid x)=\mathbb{P}(Y_{t}=y_{i}\mid X_{t}=x)=\mathds{1}_{B_{i}}(x),$
-where $\mathds{1}_{B_{i}}(x)$ is the indicator function.
-In this model, the agent has access to only the representative points $Y_{t}=y_{i}$, which indicates that the true state $X_{t}$ resides in the quantization bin $B_{i}$.
-In other words, the true state $X_{t}$ remains hidden from the agent.
-In the quantized MDP, the agent observes $Y_{t}=q(X_{t})$.
+Let $q$ be a quantizer mapping $\mathbb{X}$ to a finite set as described in Section 2.1.
 Now, consider the following Q-learning update rule for $(X_{T},U_{T})=(x,u)\in\mathbb{X}\times\mathbb{U}$:
@@
     During the exploration phase, each observation-action pair $(y,u)$ is visited infinitely often. This ensures sufficient exploration of the state-action space.
-We note that a sufficient condition for the second item above is that the state process $\{X_{t}\}_{t\geq 0}$ is positive Harris recurrent.
-The following result is proved by Kara and Yüksel [11]. It states that the algorithm in (1) converges to the optimal Q-function of the approximate belief MDP under Assumption 3.1:
-We have the following convergence result:
-Theorem 12 (Kara et al., 2023, Theorem 9 [16]).
+We note that a sufficient condition for the second item above is that the state process $\{X_{t}\}_{t\geq 0}$ is positive Harris recurrent. We have the following convergence result:
+Theorem 9 (Theorem 9 [15]).
 Under Assumption 3.1, the quantized Q-learning algorithm in 1 converges almost surely to a function $Q^{*}(y,u)$ that satisfies the following fixed-point equation for every $(y,u)\in\mathbb{Y}\times\mathbb{U}$:
@@
 $C^{*}(y,u)=\mathbb{E}_{x\sim\pi_{\gamma^{*}}|x\in B_{y}}[c(x,u)]=\int_{B_{y}}\frac{\pi_{\gamma^{*}}(dx)}{\pi_{\gamma^{*}}(B_{y})}c(x,u).$ (19)
-Observe that the above is, see e.g. [15, Theorem 2.1], equal to the almost sure limit of the empirical expression on the right hand side below:
+Observe that the above is, see e.g. [14, Theorem 2.1], equal to the almost sure limit of the empirical expression on the right hand side below:
 $C^{*}(y,u)=\lim_{N\to\infty}\frac{\sum_{k=0}^{N-1}c(X_{k},U_{k})1_{\{X_{k}\in B_{y},U_{k}=u\}}}{\sum_{k=0}^{N-1}1_{\{X_{k}\in B_{y},U_{k}=u\}}}$ (20)
@@
 $P^{*}(y^{\prime}|y,u)=\mathbb{P}_{x\sim\pi_{\gamma^{*}}|x\in B_{y}}[q(X_{t+1})=y^{\prime}|X_{t}=x,U_{t}=u]=\int_{B_{y}}\frac{\pi_{\gamma^{*}}(dx)}{\pi_{\gamma^{*}}(B_{y})}{\cal T}(B_{y^{\prime}}|x,u).$ (21)
-Likewise, by [15, Theorem 2.1], the above is the almost sure empirical limit of of the right hand side below:
+Likewise, by [14, Theorem 2.1], the above is the almost sure empirical limit of of the right hand side below:
 $P^{*}(y^{\prime}|y,u)=\lim_{N\to\infty}\frac{\sum_{k=0}^{N-1}1_{\{X_{k+1}\in B_{y^{\prime}}\}}1_{\{X_{k}\in B_{y},U_{k}=u\}}}{\sum_{k=0}^{N-1}1_{\{X_{k}\in B_{y},U_{k}=u\}}}$ (22)
-An interpretation of the above result then is that one can first obtain the approximate model given with (20-22) by forcing the data into a Markovian model for both the empirical cost estimate (21) and empirical transition kernel estimate (23), and then solve the MDP as if this Markovian empirically constructed model is the actual one, instead of running Q-learning whose limit is then optimal precisely for this learned/empirically constructed model.
+An interpretation of the above result then is that one can first obtain the approximate model given with (19-21) by forcing the data into a Markovian model for both the empirical cost estimate (20) and empirical transition kernel estimate (22), and then solve the MDP as if this empirically constructed model is the actual one, instead of running Q-learning whose limit is then optimal precisely for this learned/empirically constructed model.
 A benefit of such a model-based approach is that one can have better sample complexity bounds compared with Q-learning for certain applications, see e.g.
-[35] (see also [13, Section 5.1]).
+[34] (see also [12, Section 5.1]).
 Remark 3.1.
@@
 In contrast, within the learning context, the weighting measure $\pi_{y_{i}}$ is dictated by the exploration policy and is inherently dependent on the structure of the quantization bins $\{B_{i}\}$.
-Specifically, $\pi_{y_{i}}$ corresponds to the normalized invariant distribution over bin $B_{i}$ induced by the exploration policy; that is, it is the conditional probability measure given that the state is in $B_{i}$:
-$\pi_{y_{i}}(A)=\frac{1_{\{A\in B_{i}\}}\pi_{\gamma^{*}}(A)}{\pi_{\gamma^{*}}(B_{y})},\qquad A\in B_{i},A\in{\cal B}(\mathbb{X})$
 3.3 Error Analysis and Quantizer Design for Model Learning and Quantized Q-Learning
@@
 This limitation is significant for the analysis of non-compact spaces.
-Nevertheless, we will show that it is still possible to establish a meaningful upper bound on the expected error by employing a Lyapunov-based analysis, analogous to the approach presented in Section 2.2.
-We first start with the following result which relates the approximation error to occupation measures.
 In the previous section, we studied the approximate model where the weighting measures $\hat{\pi}^{*}_{y_{i}}$ for each quantization bin were chosen freely and $\hat{\pi}^{*}_{y_{i}}$ were chosen as Dirac measures concentrated at the $\ell_{1}$ centroids (medians) of each bin $B_{i}$.
@@
 The measures $\hat{\pi}^{*}_{y_{i}}$’s are inherently determined by the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy $\gamma^{*}$.
-We note that under weak continuity and compactness of the state space conditions, the set of invariant measures is compact in the weak topology.
-By considering this fact, we introduce the following theorem:
+We first start with the following result which relates the approximation error to occupation measures.
 Theorem 10.
-Under Assumption Assumption 2.2 and Assumption 3.1, given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}\in\mathcal{X}$:
-$\left|J^{*}_{\beta}(x_{0})-\min_{v}\{Q^{*}(x_{0},v)\}\right|$
-$\quad\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{B_{i}}\|x-x^{\prime}\|\mu_{\beta}(dx)\hat{\pi}_{y_{i}}(dx^{\prime}),$ (24)
+Under Assumption Assumption 2.2 and Assumption 3.1, given a collection of quantization bins $\{B_{i}\}_{i=1}^{M}$, we have for any initial state $x_{0}$:
+$\left|J^{*}_{\beta}(x_{0})-\min_{v}\{Q^{*}(x_{0},v)\}\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\sup_{\gamma_{s}\in\Gamma_{s}}\frac{1}{1-\beta}\sum_{i=1}^{M}\int_{B_{i}}\int_{B_{i}}\|x-x^{\prime}\|\mu_{\beta}(dx)\hat{\pi}_{y_{i}}(dx^{\prime}),$ (23)
 where, $\Gamma_{s}$ represents the set of stationary policies, $\mu_{\beta}$ is the discounted occupation measure under the policy $\gamma_{s}$ and $\hat{\pi}_{y_{i}}(dx^{\prime})$ is the normalized invariant measure obtained under the exploration policy $\gamma^{*}$.
@@
 Theorem 11.
-Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$.
+Assume that the state space $\mathds{X}\subseteq\mathbb{R}^{n}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$.
 Assume the state process $\{X_{t}\}$ satisfies the following condition:
 $\sup_{x\in\mathds{X},u\in\mathds{U}}\mathbb{E}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-\alpha V(x)+b,$ (25)
-where $V(x)=|x|^{m}$, $m>1$.
+where $V(x)=\|x\|_{1}^{m}$, $m>1$.
 Then, under Assumption 2.1 and Assumption 2.2, provided that the cost function $c$ is bounded, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathbb{X}$:
-$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (27)
+$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1/n(1-1/m)})(1-\beta)}\right),$ (26)
 where $C$ is a constant, which depends on the initial state $x_{0}$, defined as:
-$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
+$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
 Proof.
-The proof is in Appendix G. ∎
+The proof is in Appendix F. ∎
 4 Concluding Remarks
@@
 Nonetheless, asymptotic near optimality can be established under both setups with explicit convergence rates.
-In particular, under Lyapunov growth conditions, we obtained explicit upper bounds which decay to zero as the number of bins approaches infinity.
 We note that, due to relative clarity in presentation especially involving the associated Lyapunov analysis, while we studied the case with the state space being $\mathbb{R}^{n}$, the analysis can be directly generalized to any normed space by adopting the required regularity conditions on the kernels and cost.
@@
 Proof.
-Step 1: Let $\mathds{X}\subseteq\mathbb{R}$ be the state space, which we partition into $M+1$ quantization bins $\{B_{1},B_{2},\ldots,B_{M},B_{M+1}\}$.
+First consider the case with $n=1$, that is $\mathds{X}\subseteq\mathbb{R}$ Step 1: Partition $\mathds{X}$ into $M+1$ quantization bins $\{B_{1},B_{2},\ldots,B_{M},B_{M+1}\}$.
 The first $M$ bins cover a compact subset $\mathcal{K}=[-\frac{N}{2},\frac{N}{2}]\subset\mathds{X}$, and the last bin $B_{M+1}$ is an overflow bin that captures the rest of the state space outside $\mathcal{K}$.
-We apply a uniform quantizer to the compact region $\mathcal{K}$, dividing it into $M$ bins of equal length. The quantization width of each bin is:
+We apply a uniform quantizer to the compact region $\mathcal{K}$, dividing it into $M$ bins of equal length.
+The quantization width of each bin is:
 $\Delta=\frac{N}{M}.$
@@
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx),$ (30)
-where $\mu_{\beta}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$.
-Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
+where $\mu_{\beta}$ is the normalized discounted occupation measure under a stationary policy $\gamma_{s}$. Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
 $\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}^{\gamma_{s}}(\mathcal{K})={\Delta}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (31)
-Step 3: In the overflow bin $B_{M+1}$, the state space may be unbounded.
-In what follows, to simplify the analysis, we will assume that the overflow bin is always mapped to state $x=0$.
-We need to bound:
+Step 3: In the overflow bin $B_{M+1}$, the state space may be unbounded, we take that the overflow bin is always mapped to state $x=0$: Thus,
 $\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}|x-x^{\prime}|\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
@@
 where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized weighing measure over the bin $B_{M+1}$ and where the last step follows since we map the overflow bin directly to 0, i.e.
-$x^{\prime}=1$ with probability 1.
+$x^{\prime}=0$ with probability 1.
 Using Hölder’s inequality to bound the expected loss over $B_{M+1}$:
 $\int_{B_{M+1}}|x|\,\mu_{\beta}^{\gamma_{s}}(dx)=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]\leq\left(\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|^{m}]\right)^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (33)
-Step 4: In the following, we bound the moment under the measure $\mu_{\beta}^{\gamma_{s}}$. Define the process $\{M_{t}\}$ for $t\geq 0$:
+Step 4: In the following, we bound the moment of the probability measure $\mu_{\beta}^{\gamma_{s}}$. Define the process $\{M_{t}\}$ for $t\geq 0$:
 $M_{t}:=\frac{V(x_{t})}{(1-\alpha)^{t}}-\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}},$
-with $M_{0}=V(x_{0})-b=|x_{0}|^{m}-b.$
-Observe the following with respect to the filtration $\{\mathcal{F}_{t}\}$, where $\mathcal{F}_{t}=\sigma(X_{1},X_{2},\dots,X_{t})$ is the natural filtration:
+with $M_{0}=V(x_{0})-b=|x_{0}|^{m}-b.$ Observe the following with respect to the filtration $\{\mathcal{F}_{t}\}$, where $\mathcal{F}_{t}=\sigma(X_{1},X_{2},\dots,X_{t})$ is the natural filtration:
 $\mathbb{E}[M_{t+1}\mid\mathcal{F}_{t}]-M_{t}$ $=\mathbb{E}\left[\frac{V(X_{t+1})}{(1-\alpha)^{t+1}}-\sum_{k=0}^{t+1}\frac{b}{(1-\alpha)^{k}}\mid\mathcal{F}_{t}\right]-\frac{V(x_{t})}{(1-\alpha)^{t}}+\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
 $=\frac{\mathbb{E}\left[V(X_{t+1})\mid\mathcal{F}_{t}\right]}{(1-\alpha)^{t+1}}-\sum_{k=0}^{t+1}\frac{b}{(1-\alpha)^{k}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
-$\leq\frac{(1-\alpha)V(x_{t})}{(1-\alpha)^{t+1}}-\frac{b}{(1-\alpha)^{t+1}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\frac{b}{(1-\alpha)^{t+1}}$
-$=0,$
+$\leq\frac{(1-\alpha)V(x_{t})}{(1-\alpha)^{t+1}}-\frac{b}{(1-\alpha)^{t+1}}-\frac{V(x_{t})}{(1-\alpha)^{t}}+\frac{b}{(1-\alpha)^{t+1}}=0,$
 where the inequality comes from the Lyapunov condition. Hence, we showed that the process $\{M_{t}\}$ is a supermartingale for all $t\geq 0$. Then, observe that for every fixed stopping time $t$:
@@
 $\mathbb{E}[V(X_{t})]$ $\leq(V(X_{0})-b)(1-\alpha)^{t}+(1-\alpha)^{t}\sum_{k=0}^{t}\frac{b}{(1-\alpha)^{k}}$
-$=(V(X_{0})-b)(1-\alpha)^{t}+b\sum_{k=0}^{t}(1-\alpha)^{k}$
-$=(V(X_{0})-b)(1-\alpha)^{t}+b\cdot\frac{1-(1-\alpha)^{t+1}}{\alpha}.$
+$=(V(X_{0})-b)(1-\alpha)^{t}+b\sum_{k=0}^{t}(1-\alpha)^{k}=(V(X_{0})-b)(1-\alpha)^{t}+b\cdot\frac{1-(1-\alpha)^{t+1}}{\alpha}.$
 Note that
@@
 $=\frac{V(x_{0})(1-\beta)+b\beta}{1-\beta(1-\alpha)}=:C,$ (34)
-which shows that the error bound $C$ depends on the initial state $x_{0}$ and other constants.
 Step 5: Returning back to equation (33), we find:
@@
 $\leq C^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$ (35)
-where $C$ is defined in (35).
-Combining the bounds from (32), (C) and (36), the total expected loss is bounded by:
+where $C$ is defined in (34). Combining the bounds from (31), (C) and (35):
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq{\Delta}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+C^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (36)
@@
 where $M$ is the number of quantization bins, $m>1$ is the moment in the Lyapunov function, and $C$ is (defined in (34)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$.
-Appendix D Proof of Theorem 8
-Proof.
-Step 1: Let the state space be $\mathds{X}\subseteq\mathbb{R}^{n}$ and fix a side–length parameter $N>0$. Choose an integer $k\geq 1$ and set
+This completes the case with $n=1$. We now consider the case where $\mathds{X}\subseteq\mathbb{R}^{n}$ with $n\geq 1$.
+Now, fix a side length parameter $N>0$. Choose an integer $k\geq 1$ and set
 $M:=k^{n}\qquad\bigl(\text{so that }k=M^{1/n}\bigr).$
@@
 $B_{M+1}\;:=\mathds{X}\setminus\mathcal{K},$
-which is the set of states outside the compact (granular) grid.
-Step 2: For any state $x\in\mathcal{K}$, due to the uniform quantization, we have that
-$L(x)\leq{n\Delta}$
-We decompose the expected loss into two parts as in the proof of Theorem 7:
+which is the set of states outside the compact (granular) grid. We decompose the expected loss into two parts as in the proof of Theorem 6:
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\mu_{\beta}^{\gamma_{s}}(dx),$ (39)
@@
 $\int_{\mathcal{K}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)\leq{n\,\Delta\,\mu_{\beta}^{\gamma_{s}}\!(\mathcal{K})}={n\,\Delta\bigl[1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\bigr]}.$ (40)
-Step 3 : Inside the overflow bin $B_{M+1}\subseteq\mathds{X}$ the state space is unbounded.
-In the quantizer, for simplicity in analysis, we may take that the overflow bin is directly mapped to state $x=0$.
-We must bound:
+Inside the overflow bin $B_{M+1}\subseteq\mathds{X}$, we take again that the overflow bin is directly mapped to state $x=0$ and thus,
 $\int_{B_{M+1}}L(x)\mu_{\beta}^{\gamma_{s}}(dx)=\int_{B_{M+1}}\int_{B_{M+1}}\|x-x^{\prime}\|_{1}\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
-$\leq\int_{B_{M+1}}\int_{B_{M+1}}\bigl(\|x\|_{1}+\|x^{\prime}\|_{1}\bigr)\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$
-$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]$
+$\leq\int_{B_{M+1}}\int_{B_{M+1}}\bigl(\|x\|_{1}+\|x^{\prime}\|_{1}\bigr)\hat{\pi}^{*}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]$
 where the last step follows since we assume that $x^{\prime}=0$ always for the overflow bin. We apply Hölder’s inequality with exponent $m>1$:
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}\mathbf{1}_{B_{M+1}}(X)\bigr]\leq\Bigl(\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}\bigl[\|X\|_{1}^{m}\bigr]\Bigr)^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (41)
-Step 4: Repeating the arguments in Step 4 of the proof of Theorem 7, we can show that
+Repeating the arguments in Step 4,
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]\leq\frac{V(x_{0})(1-\beta)+b\beta}{1-\beta(1-\alpha)}=:C$
@@
 $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq{n\Delta\bigl[1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\bigr]}+C^{1/m}\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m}.$ (42)
-We select $N\;=\;2\bigl(C\,M\bigr)^{1/m}.$ By Markov’s Inequality:
+We select $N=2\bigl(C\,k\bigr)^{1/m}.$ By Markov’s Inequality:
 $\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $={\mu_{\beta}^{\gamma_{s}}(\{x={x_{1},\cdots,x_{n}}:\min_{i=1,\cdots,n}|x_{i}|\geq N/2\})\leq}\mathop{\rm Pr}\bigl[\|X\|_{1}\geq N/2\bigr]$
-$\leq\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]}{(\tfrac{N}{2})^{m}}=\frac{C}{C\,M}=\frac{1}{M}.$
+$\leq\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}^{m}]}{(\tfrac{N}{2})^{m}}=\frac{C}{C\,k}=\frac{1}{k}.$
 With $k=M^{1/n}$ interior cells per axis, the hyper-cube side length is
-$\Delta=\frac{N}{k}=2C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}.$
-Inserting these bounds into (43) and assuming that $m>n$:
-$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}\left(\frac{1}{M}\right)^{1-\frac{1}{m}}$
-$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{-\left(1-\frac{1}{m}\right)}$
-$\leq 2nC^{1/m}M^{\frac{1}{m}-\frac{1}{n}}+C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
-$=(2n+1)C^{1/m}M^{\frac{1}{m}-\frac{1}{n}}$
-$=(2n+1)C^{1/m}M^{-\left(\frac{1}{n}-\frac{1}{m}\right)}.$
-As a result, as $M\to\infty$, the total expected loss $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$ since $m>n$.
-Under necessary conditions, we obtain:
-$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{1/n-1/m})(1-\beta)},$ (44)
-where $n$ is the dimension of the state space, $M$ is the number of quantization bins, $m>n$ is the moment in the Lyapunov function, and $C$ is (defined in (35)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$.
+$\Delta=\frac{N}{k}=2C^{\frac{1}{m}}k^{(1-\frac{1}{m})}.$
+Inserting these bounds into (42):
+$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\leq 2nC^{1/m}\left(\frac{1}{k}\right)^{1-\frac{1}{m}}+C^{1/m}\left(\frac{1}{k}\right)^{1-\frac{1}{m}}\leq(2n+1)C^{1/m}k^{-(1-\frac{1}{m})}$
+As $M\to\infty$, and so as $k\to\infty$, we have $\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]\to 0$. Thus,
+$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{\frac{1}{n}(1-\frac{1}{m})})(1-\beta)},$ (43)
+where $n$ is the dimension of the state space, $M$ is the number of quantization bins, $m$ is the moment in the Lyapunov function, and $C$ is (defined in (34)) the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}^{\gamma_{s}}$.
 ∎
-Appendix E Proof of Theorem 9
+Appendix D Proof of Theorem 7
 Proof.
@@
 $\hat{h}(y)=\inf_{u\in\mathds{U}}\left\{C^{*}(y,u)+\sum_{y_{1}}\hat{h}(y_{1})P^{*}(y_{1}|y,u)\right\}-\int_{\mathds{X}}\hat{h}(q(x))\mu(dx).$ (45)
+According to Verification Theorem for average cost criterion (see [9] and [1]), the long-term average cost $J^{*}(x)$ for the original model and $\hat{J}(x)$ for the discretized model are given by:
+$J^{*}(x)=j^{*}=\int_{\mathds{X}}h(x)\mu(dx),$
+$\hat{J}(x)=\hat{j}=\int_{\mathds{X}}\hat{h}(x)\mu(dx),$
+[…]
 For the rest of the proof, we simply use the same notation for $C^{*}$, and $\hat{h}$ by extending them as constant over the quantization bins, that is we override the notation and use $\hat{h}(x)$ for $\hat{h}(q(x))$ and $C^{*}(x,u)$ for $C^{*}(q(x),u)$.
@@
 $\hat{h}(x)=\inf_{u\in\mathds{U}}\left\{C^{*}(x,u)+\int\hat{h}(x_{1})\hat{\mathcal{T}}(dx_{1}|x,u)\right\}-\int_{\mathds{X}}\hat{h}(x)\mu(dx).$ (46)
-According to Verification Theorem for average cost criterion (see [9] and [1]), the long-term average cost $J^{*}(x)$ for the original model and $\hat{J}(x)$ for the discretized model are given by:
-$J^{*}(x)=j^{*}=\int_{\mathds{X}}h(x)\mu(dx),$
-$\hat{J}(x)=\hat{j}=\int_{\mathds{X}}\hat{h}(x)\mu(dx),$
-[…]
 We denote by $\mathcal{T}^{-}(\cdot|x,u):=\mathcal{T}(\cdot|x,u)-\mu(\cdot)$ and $\hat{\mathcal{T}}^{-}(\cdot|x,u):=\hat{\mathcal{T}}(\cdot|x,u)-\mu(\cdot)$. Note that
@@
 $h_{k}(x)\leq\|c\|_{\infty}+\sup_{u}\left\{\int h_{k-1}(x_{1})\hat{\mathcal{T}}^{-}(dx_{1}|x,u)\right\}\leq\|c\|_{\infty}+\|h_{k-1}\|_{\infty}(1-\mu(\mathds{X})).$
-Letting $\alpha:=(1-\mu(\mathds{X}))$, and repeating this step, we write
-$\|h_{k}\|_{\infty}\leq\|c\|_{\infty}\sum_{t=0}^{k-1}\alpha^{t}.$
-Finally, using the fact that the operator $\hat{T}$ is a contraction under the supremum norm under Assumption 2.3 with the fixed point $\hat{h}$, we conclude
-$\|\hat{h}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\alpha}=\frac{\|c\|_{\infty}}{\mu(\mathds{X})}.$
-Identical steps can be used to show that $\|h\|_{\infty}\leq\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$. Combining what we have so far, we write
-[…]
-$\leq\left(\alpha_{c}+\frac{\|c\|_{\infty}\alpha_{T}}{\mu(\mathds{X})}\right)\int L(x)\pi_{\gamma_{s}}(dx)$
-which completes the proof.
-Appendix F Proof of Theorem 11
+Letting $\alpha:=(1-\mu(\mathds{X}))$, and repeating this step, we write $\|h_{k}\|_{\infty}\leq\|c\|_{\infty}\sum_{t=0}^{k-1}\alpha^{t}$.
+Finally, using the fact that the operator $\hat{T}$ is a contraction under the supremum norm under Assumption 2.3 with the fixed point $\hat{h}$, we conclude $\|\hat{h}\|_{\infty}\leq\frac{\|c\|_{\infty}}{1-\alpha}=\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$.
+Identical steps can be used to show that $\|h\|_{\infty}\leq\frac{\|c\|_{\infty}}{\mu(\mathds{X})}$.
+Combining what we have so far, we write
+$|j^{*}-\hat{j}|=\left|\int h(x)\mu(dx)-\int\hat{h}(x)\mu(dx)\right|\leq\int V(x)\mu(dx)\leq\left(\alpha_{c}+\frac{\|c\|_{\infty}\alpha_{T}}{\mu(\mathds{X})}\right)\int L(x)\pi_{\gamma_{s}}(dx).$
+Appendix E Proof of Theorem 8
 Proof.
-First we assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $V:\mathds{X}\to[0,\infty)$, $f:\mathds{X}\to[\epsilon,\infty)$ for some $\epsilon>0$.
-Assume the state process $\{X_{t}\}$ satisfies the following condition:
-$\sup_{x\in\mathds{X},u\in\mathds{U}}\mathbb{E}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$ (48)
-where $f(x)=|x|^{m}$, $m>1$.
-Then, we choose a collection of quantization bin $\{B_{i}\}_{i}^{M+1}$ such that the first $M$ bins quantize the compact set $(bM)^{1/m},(bM)^{1/m}]$ uniformly and the last bin $B_{M+1}$ is the overflow bin that captures the rest of the state space.
+Choose a collection of quantization bin $\{B_{i}\}_{i}^{M+1}$ such that the first $M$ bins quantize the compact set $(bk)^{1/m},(bk)^{1/m}]^{n}$ uniformly where $M=k^{n}$ and the last bin $B_{M+1}$ is the overflow bin that captures the rest of the state space.
 Then, the total expected distortion becomes:
 $\mathbb{E}_{\pi_{\gamma_{s}}}[L(X)]=\int_{\mathcal{K}}L(x)\,\pi_{\gamma_{s}}(dx)+\int_{B_{M+1}}L(x)\,\pi_{\gamma_{s}}(dx)$
-In the overflow bin $B_{M+1}$, the state space may be unbounded, and the quantization error $L(x)=\|x-y_{M+1}\|$ can be large. We need to bound:
-$\int_{B_{M+1}}L(x)\,\pi_{\gamma_{s}}(dx)=\int_{B_{M+1}}|x-y_{M+1}|\,\pi_{\gamma_{s}}(dx),$ (49)
-where $y_{M}$ is the representative point for bin $B_{M+1}$.
-We choose $y_{M+1}=0$ (or any fixed point), so that $L(x)=|x|$ for $x\in B_{M+1}$.
-By our assumption, there exists a Lyapunov function $V$ that satisfies inequality (48). Then, by Theorem 10, for any initial state $x_{0}=x$, we have:
-$\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq V(x)+\mathbb{E}\left[\sum_{t=0}^{T-1}b\right]=V(x)+bT,$
-where $T$ is a deterministic stopping time. By taking the limit for both sides, we get:
-$\limsup_{T\to\infty}\frac{1}{T}\mathbb{E}_{x}\left[\sum_{t=0}^{T-1}f(x_{t})\right]\leq\limsup_{T\to\infty}\frac{1}{T}(V(x)+bT)=b,$
-Now, consider any invariant probability measure $\pi$. Recall that we assigned $f=|x|^{m},m>1$. Fix $N>0$ and define $f_{N}(x)=\min(|x|^{m},N)$. Then, we can write (see [34, p. 75]):
-$\int_{\mathds{X}}f_{N}(x)\pi(dx)$ $=\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{x_{0}\sim\pi}\left[E_{x_{0}}[f_{N}(X_{t})]\right]$
-$\leq\mathbb{E}_{\pi}\left[\limsup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}E_{x_{0}}[f_{N}(X_{t})]\right]\leq b,$
-where the first inequality comes from applying Fatou’s Lemma since $f_{N}(x)$ is bounded by $N$.
-By Taking the limit $N\to\infty$ and noting that $f_{N}(x)\to|x|^{m}$ pointwise and monotonically, by Monotone Convergence Theorem, we obtain $\int_{\mathds{X}}|x|^{m}\pi(dx)\leq b$.
-Finally, since this holds for any invariant probability measure $\pi$ and $|x|^{m}$ is non-negative, we conclude that $\mathbb{E}[|X_{t}|^{m}]\leq b,\forall t\geq 0.$ By following the same procedure in the proof of Theorem 7, we obtain the following bound:
+As earlier, we choose $y_{M+1}=0$ (or any fixed point), so that $L(x)=\|x\|_{1}$ for $x\in B_{M+1}$.
+By our assumption, there exists a Lyapunov function $V$ that satisfies inequality (11).
+Thus, by [33, Theorem 4.2.5] (which builds critically on the Comparison Theorem [18, Theorem 14.2.2]), under any invariant probability measure $\pi$: $\int_{\mathds{X}}\|x\|_{1}^{m}\pi(dx)\leq b$
+By following the same procedure as in the proof of Theorem 6, we obtain the following bound:
 $\mathbb{E}_{\pi_{\gamma_{s}}}[L(X)]$ $\leq\Delta\cdot\left(1-\pi_{\gamma_{s}}(B_{M+1})\right)+b^{1/m}\cdot\pi_{\gamma_{s}}(B_{M+1})^{1-1/m}$ (47)
-$\leq\Delta\ +\frac{b^{1/m}}{M^{1-1/m}}\leq\frac{2b^{1/m}}{M^{1-1/m}}+\frac{b^{1/m}}{M^{1-1/m}}=\frac{3b^{1/m}}{M^{1-1/m}}.$ (51)
+$\leq\Delta\ +\frac{b^{1/m}}{k^{1-1/m}}\leq\frac{2nb^{1/m}}{k^{1-1/m}}+\frac{b^{1/m}}{k^{1-1/m}}=\frac{(2n+1)b^{1/m}}{k^{1-1/m}}.$ (48)
 By combining this bound with Theorem 7 under the given assumptions, we obtain:
-$\left|\hat{J}_{avg}(x_{0})-J^{*}_{avg}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{(M^{1-1/m})},$ (52)
+$\left|\hat{J}_{avg}(x_{0})-J^{*}_{avg}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{(2n+1)b^{1/m}}{(M^{1/n(1-1/m)})},$ (49)
 where $M$ is the number of quantization bins, $m>1$ is the moment in Lyapunov function and $b$ is the uniform bound on the $m$-th moment of $X_{t}$.
@@
 ∎
-Appendix G Proof of Theorem 14
+Appendix F Proof of Theorem 11
 Proof.
-Step1: We partition the state space similar to the quantization in Theorem 7 by choosing the compact set $\mathcal{K}=[-(CM)^{1/m},(CM)^{1/m}]$.
+Step1: We partition the state space similar to the quantization in Theorem 6 by choosing the compact set $\mathcal{K}=[-(Ck)^{1/m},(Ck)^{1/m}]$ with $k^{n}=M$, ($k=M^{1/n}$).
 Then, for any state $x\in\mathcal{K}$, the quantization error $L(x)$ satisfies:
-$L(x)\leq{\Delta}=\frac{2(CM)^{1/m}}{M}=\frac{2C^{1/m}}{M^{1-1/m}}.$ (53)
+$L(x)\leq{\Delta}=\frac{2n(Ck)^{1/m}}{k}=\frac{2nC^{1/m}}{k^{1-1/m}}=\frac{2nC^{1/m}}{M^{1/n(1-1/m)}}.$ (50)
 We decompose the total expected loss into two parts:
@@
 Step 2: Since $L(x)\leq{\Delta}$ for $x\in\mathcal{K}$, we have:
-$\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}(\mathcal{K})=\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (55)
+$\int_{\mathcal{K}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)\leq{\Delta}\cdot\mu_{\beta}(\mathcal{K})=\frac{2nC^{1/m}}{M^{1/n(1-1/m)}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right).$ (52)
 Step 3 (Two parallel arguments): In the overflow bin $B_{M+1}$, the state space may be unbounded.
@@
 Thus, the bound becomes:
-$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}|x-x^{\prime}|\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (56)
-$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(|x|+|x^{\prime}|\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (57)
-$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]+\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1}),$ (58)
+$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $=\int_{B_{M+1}}\int_{B_{M+1}}\|x-x^{\prime}\|_{1}\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (53)
+$\leq\int_{B_{M+1}}\int_{B_{M+1}}\left(\|x\|_{1}+\|x^{\prime}\|_{1}\right)\hat{\pi}_{M+1}(dx^{\prime})\mu_{\beta}^{\gamma_{s}}(dx)$ (54)
+$=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]+\mathbb{E}_{\hat{\pi}_{M+1}}[\|X^{\prime}\|_{1}]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1}),$ (55)
 where $\hat{\pi}^{*}_{M+1}(dx^{\prime})$ is the normalized invariant measure of the state process over the bin $B_{M+1}$ under the exploration policy. The second term is handled in two ways:
 3.a
-Observe that if $x_{0}\sim\pi_{\gamma^{*}}$ then $\mu_{\beta}^{\gamma_{s}}=\pi_{\gamma^{*}}$ and therefore the terms in the summation above, that is,
-$\mathbb{E}_{\hat{\pi}^{*}_{M+1}}[|X^{\prime}|]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mu_{\beta}^{\gamma_{s}}(B_{M+1})\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]$
-will be identical.
+Observe that if $x_{0}\sim\pi_{\gamma^{*}}$ then $\mu_{\beta}^{\gamma_{s}}=\pi_{\gamma^{*}}$ and therefore the terms in the summation above will be identical, that is,
+$\mathbb{E}_{\hat{\pi}^{*}_{M+1}}[\|X^{\prime}\|_{1}]\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mu_{\beta}^{\gamma_{s}}(B_{M+1})\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)].$
 Similar to the proof of Theorem 6, by the law of the iterated expectations, we can show that when the initial state is not fixed but given by $x_{0}\sim\pi_{\gamma^{*}}$:
-$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|X|\mathds{1}_{B_{M+1}}(X)]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\mathbb{E}[|X|\mathds{1}_{B_{M+1}}(X)|X_{0}=x]]\leq\hat{C}^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$
+$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|X\|\mathds{1}_{B_{M+1}}(X)]=\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\mathbb{E}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)|X_{0}=x]]\leq\hat{C}^{1/m}\cdot\mu_{\beta}^{\gamma_{s}}(B_{M+1})^{1-1/m},$
 where
-$\hat{C}:=\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[|x_{0}|]^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
+$\hat{C}:=\frac{\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[\|x_{0}\|_{1}]^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
 By using Markov’s inequality, we obtain:
-$\mu_{\beta}^{\gamma_{s}}(B_{M+1})=\mathbb{P}\left(|X|\geq(CM)^{1/m}\right)\leq\frac{\mathbb{E}[|X|^{m}]}{\left((CM)^{1/m}\right)^{m}}\leq\frac{1}{M},$ (59)
+$\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $={\mu_{\beta}^{\gamma_{s}}(\{x={x_{1},\cdots,x_{n}}:\min_{i=1,\cdots,n}|x_{i}|\geq(Ck)^{1/m}\})\leq}\mathop{\rm Pr}\bigl[\|X\|_{1}\geq(Ck)^{1/m}\bigr]\leq\frac{1}{k}$ (56)
 which gives:
-$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2\hat{C}^{1/m}}{M^{1-1/m}}$ (60)
+$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2\hat{C}^{1/m}}{k^{1-1/m}}=\frac{2\hat{C}^{1/m}}{M^{1/n(1-1/m)}}$ (57)
 3.b
 We know that $\hat{\pi}_{M+1}$ is the normalized measure of the invariant distribution $\pi_{\gamma^{*}}$ under the exploration policy. Then, we have:
-$\mathbb{E}_{\hat{\pi}_{M+1}}[|X|]=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}\cdot\pi_{\gamma^{*}}(B_{M+1})^{1-1/m}}{\pi_{\gamma^{*}}(B_{M+1})}=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}$
-By using Comparison Theorem 10 and using a similar procedure in Section 2.3, under the Lyapunov condition in (26) we obtain:
-$\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]\leq\frac{b}{\alpha}.$ (61)
-For the overflow set $B_{M+1}=\{|x|\geq(CM)^{1/m}\}$ we have by Markov’s inequality:
-$\pi_{\gamma^{*}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]}{(CM)^{m/m}}\leq\frac{b/\alpha}{CM}.$
+$\mathbb{E}_{\hat{\pi}_{M+1}}[\|X\|_{1}]=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}\mathds{1}_{B_{M+1}}(X)]}{\pi_{\gamma^{*}}(B_{M+1})}\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|^{m}]^{1/m}\cdot\pi_{\gamma^{*}}(B_{M+1})^{1-1/m}}{\pi_{\gamma^{*}}(B_{M+1})}=\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}$
+By [18, Theorem 14.2.2], under the Lyapunov condition in (25) we obtain:
+$\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]\leq\frac{b}{\alpha}.$ (58)
+For the overflow set $B_{M+1}$ we have by Markov’s inequality:
+$\pi_{\gamma^{*}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]}{(Ck)^{m/m}}\leq\frac{b/\alpha}{Ck}.$
 Putting the pieces together:
-$\mathbb{E}_{\hat{\pi}_{M+1}}[|X^{\prime}|]\mu_{\beta}^{\gamma_{s}}(B_{M+1})$ $\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[|X|^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}\cdot\frac{1}{M}$
-$=\Bigl(\tfrac{b}{\alpha}\Bigr)^{1/m}\Bigl(\tfrac{C}{b/\alpha}\Bigr)^{1/m}M^{-(1-1/m)}=\frac{C^{1/m}}{M^{1-1/m}}$
+$\mathbb{E}_{\hat{\pi}_{M+1}}[\|X^{\prime}\|_{1}]\mu_{\beta}^{\gamma_{s}}(B_{M+1})\leq\frac{\mathbb{E}_{\pi_{\gamma^{*}}}[\|X\|_{1}^{m}]^{1/m}}{\pi_{\gamma^{*}}(B_{M+1})^{1/m}}\cdot\frac{1}{k}=\Bigl(\tfrac{b}{\alpha}\Bigr)^{1/m}\Bigl(\tfrac{C}{b/\alpha}\Bigr)^{1/m}k^{-(1-1/m)}=\frac{C^{1/m}}{M^{1/n(1-1/m)}}$
 and
-$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2C^{1/m}}{M^{1-1/m}},$
+$\int_{B_{M+1}}L(x)\,\mu_{\beta}^{\gamma_{s}}(dx)$ $\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}},$
 which goes to zero as M goes to infinity. Thus the expected quantization error for the overflow bin vanishes as the number of quantization bin, M, increases.
@@
 Step 4: Now, combining the bounds we obtained so far, we get:
-$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq\frac{2C^{1/m}}{M^{1-1/m}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+\frac{2C^{1/m}}{M^{1-1/m}}$
-$\leq\frac{2C^{1/m}}{M^{1-1/m}}+\frac{2C^{1/m}}{M^{1-1/m}}=\frac{4C^{1/m}}{M^{1-1/m}}.$ (62)
+$\mathbb{E}_{\mu_{\beta}^{\gamma_{s}}}[L(X)]$ $\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}}\cdot\left(1-\mu_{\beta}^{\gamma_{s}}(B_{M+1})\right)+\frac{2C^{1/m}}{M^{1/n(1-1/m)}}$
+$\leq\frac{2C^{1/m}}{M^{1/n(1-1/m)}}+\frac{2C^{1/m}}{M^{1/n(1-1/m)}}=\frac{4C^{1/m}}{M^{1/n(1-1/m)}}.$ (59)
 Obviously, $m>1$ implies 11/m>01-1/m>0.
 Step 5: Let By combining this bound with Theorem 3 under the necessary assumptions, we obtain:
-$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1-1/m})(1-\beta)}\right),$ (63)
+$\left|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\right|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\left(\frac{4C^{1/m}}{(M^{1/n(1-1/m)})(1-\beta)}\right),$ (60)
 where $M$ is the number of quantization bins, $m>1$ is the moment in Lyapunov function, $C$ is the uniform bound on the $m$-th moment of $X_{t}$ with respect to measure $\mu_{\beta}$. ∎